### v2

# Average-Case Deterministic Query Complexity of Boolean Functions with Fixed Weight

Abstract
We explore the average-case deterministic query complexity of boolean functions under a uniform distribution, denoted by $\mathrm{D}_{\mathrm{ave}}(f)$, the minimum average depth of zero-error decision tree computing a boolean function $f$. This measure has found several applications across diverse fields, yet its understanding is limited. We study $\mathrm{D}_{\mathrm{ave}}(f)$ of several functions, including the penalty shoot-out function, symmetric functions, linear threshold functions and the tribes functions.
We prove $\mathrm{D}_{\mathrm{ave}}(f)\leq\max\{\log\frac{\operatorname{wt}(f)}{\log n}+O(\log\log\frac{\operatorname{wt}(f)}{\log n}),O(1)\}$ for every $n$-variable boolean function $f$, where $\operatorname{wt}(f)$ denotes the weight (the number of inputs on which $f$ outputs $1$). For any $4\log n\leq m(n)\leq 2^{n-1}$, we prove the upper bound is tight up to an additive logarithmic term for almost all $n$-variable boolean functions with weight $\operatorname{wt}(f)=m(n)$.
Using Håstad’s switching lemma or Rossman’s switching lemma [Comput. Complexity Conf. 137, 2019], one can derive $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{O(w)})$ or $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{O(\log s)})$ for CNF/DNF formulas of width $w$ or size $s$, respectively. We show that, for any $w\geq\log n+\log\log n+3$, there exists a DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$ such that $\mathrm{D}_{\mathrm{ave}}(f)=n(1-\frac{\log n}{\Theta(w)})$. In other words, we show the criticality upper bounds $O(w)$ and $O(\log s)$ are tight up to a multiplicative $\log n$ factor, providing evidence on the tightness of the switching lemmas.
Keywords— average-case query complexity, decision tree, weight, criticality, switching lemma
1 Introduction
The average-case deterministic query complexity of a boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ under a uniform distribution [^1] In [AD01], the average-case query complexity under distribution $\mu$ is denoted by $\mathrm{D}^{\mu}(f)$; throughout this paper, we always assume the input distribution is uniform., denoted by $\mathrm{D}_{\mathrm{ave}}(f)$, is the minimum average depth of zero-error decision trees computing $f$. This notion serves as a natural average-case analogy of the classic deterministic query complexity and has found applications in query complexity, boolean function analysis, learning algorithms, game theory, circuit complexity and percolation theory. Besides that, $\mathrm{D}_{\mathrm{ave}}(f)$ falls outside the class of polynomially-related measures including $\mathrm{D}(f)$, $\mathrm{R}(f)$, $\mathrm{C}(f)$, $\mathrm{bs}(f)$, $\mathrm{s}(f)$ (see the summaries in [BD02, ABK16, ABK+21] and Huang’s proof of the Sensitivity Conjecture [Hua19].); $\mathrm{D}_{\mathrm{ave}}(f)$ is a measure with limited understanding. In addition, this work is also inspired by Rossman’s average-case circuit lower bounds of detecting $k$-clique on Erdős–Rényi graphs [Ros08, Ros14]. In this paper, we aim to initiate a comprehensive study on $\mathrm{D}_{\mathrm{ave}}(f)$, exploring its implications and applications.
1.1 Background
To our knowledge, Ambainis and de Wolf were the first to introduce the concept of average-case query complexity. They showed super-polynomial gaps between average-case deterministic query complexity, average-case bounded-error randomized query complexity, and average-case quantum query complexity [AD01]. In another paper, Ambainis et al. [Amb+16] studied the worst-case quantum query complexity for fixed-weight functions, and proved that almost all functions (with a fixed number of variables and fixed weight) have the same quantum query complexity (up to a constant factor).
O’Donnell, Saks, Schramm, and Servedio established the OSSS inequality [OSS+05] and the OS inequality [OS06], revealing a connection between average-case deterministic query complexity, influences, average sensitivity, and Fourier coefficients. As a result, any balanced function with $\mathrm{D}(f)\leq d$ has a variable with influence at least $1/d$. Refer to [Lee10, JZ11] for simpler proofs of the OSSS inequality. O’Donnell and Servedio [OS06] discovered the OS inequality while studying learning algorithms; the OS inequality serves as a key ingredient to learning any monotone boolean function, under the uniform distribution, in time polynomial in $n$ and the decision tree size of $f$.
In the early days of theoretical computer science, Yao [Yao77] noticed that the average-case deterministic query complexity $\mathrm{D}_{\mathrm{ave}}^{\mu}(f)$ (with respect to any distribution $\mu$) lower bounds the zero-error randomized query complexity, that is, $\mathrm{D}_{\mathrm{ave}}(f)\leq\mathrm{R}_{0}(f)$. (Moreover, Yao’s min-max principle says the maximum $\mathrm{D}_{\mathrm{ave}}^{\mu}(f)$, ranging over all distributions $\mu$, coincides with the zero-error randomized query complexity $\mathrm{R}_{0}(f)$.) For example, the zero-error randomized decision tree complexity of the recursive ternary majority function $\mathrm{Maj}_{3}^{\otimes d}$, which is bounded below by $\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{3}^{\otimes d})$, has been extensively studied [SW86, JKS03, Leo13, Mag+16]; the notion of average-case deterministic query complexity sometimes plays a role in the lower bound proof. Now the best known bound is $\Omega(2.57143^{d})\leq\mathrm{R}_{0}(\mathrm{Maj}_{3}^{\otimes d})\leq O(2.64944^{d})$ [Mag+16]. In addition, when $\mu$ is uniform, $\mathrm{D}_{\mathrm{ave}}^{\mu}(\mathrm{Maj}_{3}^{\otimes d})=\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{3}^{\otimes d})=O(2.5^{d})$ (Exercise 8.38 in [ODo14]).
By applying the OSSS and the OS inequalities, O’Donnell et al. [OSS+05] proved that $\mathrm{R}_{0}(f)\geq\mathrm{D}_{\mathrm{ave}}^{(p)}(f)\geq(n/\sqrt{4p(1-p)})^{2/3}$ for any nontrivial monotone graph property $f$ with critical probability $p$. Here, $\mathrm{D}_{\mathrm{ave}}^{(p)}(f)$ is defined under the $p$-biased distribution. The result of O’Donnell et al. made progress on Yao’s conjecture [Yao77]: $\mathrm{R}_{0}(f)=\Omega(n)$ for every nontrivial monotone graph property; on the other hand, when $p=1/2$, Benjamini et al. proved that the lower bound $\mathrm{D}_{\mathrm{ave}}(f)\geq\Omega(n^{2/3})$ is tight up to a multiplicative logarithmic factor [BSW05].
The most surprising connection (application) arose in game theory. Peres, Schramm, Sheffield, and Wilson [PSS+07] studied the random-turn HEX game, in which players determine who plays next by tossing a coin before each round. They proved that the expected playing time (under optimal play) coincides with $\mathrm{D}_{\mathrm{ave}}(f)$, where $f$ is the $L\times L$ hexagonal cells connectivity function. Using the OS inequality and the results of Smirnov and Werner on percolation [SW01], Peres et al. proved a lower bound $L^{1.5+o(1)}$ on the expected playing time on an $L\times L$ board.
1.2 Our results
In query complexity, the worst-case measures $\mathrm{D}(f)$, $\mathrm{R}(f)$, $\mathrm{C}(f)$, $\mathrm{bs}(f)$, $\mathrm{s}(f)$ are all polynomially related. In contrast, we construct monotone balanced functions such that the polynomial relation between any two of the average-case measures $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$, $\mathrm{C}_{\mathrm{ave}}(f)$, $\mathrm{bs}_{\mathrm{ave}}(f)$ and $\mathrm{s}_{\mathrm{ave}}(f)$ does not hold. (It was already known that $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$ and $\mathrm{Q}_{\mathrm{ave}}(f)$ are not polynomially related [AD01].)
As complexity measures, $\mathrm{D}_{\mathrm{ave}}(f)$ and $\mathrm{D}(f)$ behave differently. Altering the value of $f(x)$ for a single input $x\in\{0,1\}^{n}$ may result in an $\Omega(n)$ change of $\mathrm{D}(f)$, since $\mathrm{D}(f)=n$ as long as $\operatorname{wt}(f)$ is odd. On the other hand, we prove $\mathrm{D}_{\mathrm{ave}}(f)\leq\max\{O(\log\operatorname{wt}(f)-\log\log n),O(1)\}$ for every $n$-variable boolean function $f$ (Theorem 1.1). As an application, $|\mathrm{D}_{\mathrm{ave}}(f)-\mathrm{D}_{\mathrm{ave}}(g)|<O(\log\operatorname{dist}(f,g))$; in other words, $\mathrm{D}_{\mathrm{ave}}(f)$ is Lipschitz continuous, after normalizing the Hamming distance between boolean functions. In addition, Theorem 1.1 implies $|\mathrm{D}_{\mathrm{ave}}(f)-\mathrm{D}_{\mathrm{ave}}(g)|<O(1)$ whenever $f,g$ disagree on $O(\log n)$ points at most, i.e., the Hamming distance $\operatorname{dist}(f,g)=O(\log n)$.
Put differently, Theorem 1.1 reveals a connection between $\mathrm{D}_{\mathrm{ave}}(f)$ and the zero-order Fourier coefficient $\widehat{f}(\{\emptyset\})$ (or equivalently $\mathbb{E}[f]$). Combining our upper bound with the FKN theorem [FKN02] and the KS theorem [KS03], one can show that boolean functions with Fourier weights concentrated on the low levels have small $\mathrm{D}_{\mathrm{ave}}(f)$.
Theorem 1.1.
For every boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$, if the weight $\operatorname{wt}(f)\geq 4\log n$, then
$\mathrm{D}_{\mathrm{ave}}(f)\leq\log\frac{\operatorname{wt}(f)}{\log n}+O\left(\log\log\frac{\operatorname{wt}(f)}{\log n}\right).$ (1)
Otherwise, $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$.
Next, we prove Theorem 1.1 is tight up to an additive logarithmic term for almost all fixed-weight functions. Ambainis et al. [Amb+16] showed a similar phenomenon for quantum query complexity. Besides that, Theorem 1.2 enables us to implicitly select a boolean function with fixed weight and large $\mathrm{D}_{\mathrm{ave}}(f)$, serving an ingredient in the proof of our next result (Theorem 1.3).
Theorem 1.2.
Let $m:\mathbb{N}\to\mathbb{N}$ be a function such that $4\log n\leq m(n)\leq 2^{n-1}$. For almost all boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$ with the same weight $\operatorname{wt}(f)=m(n)$,
$\mathrm{D}_{\mathrm{ave}}(f)\geq\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)\geq\log\frac{\operatorname{wt}(f)}{\log n}-O\left(\log\log\frac{\operatorname{wt}(f)}{\log n}\right),$ (2)
where $\mathrm{C}_{x}(f)$ denotes the size of the smallest certificate on input $x$.
Note that, for boolean functions with $\operatorname{wt}(f)\geq 2^{n-1}$, we have similar bounds by replacing $\operatorname{wt}(f)$ by $2^{n}-\operatorname{wt}(f)$ in (1) and (2), since $\mathrm{D}_{\mathrm{ave}}(f)=\mathrm{D}_{\mathrm{ave}}(\neg f)$ for any $f$, where $\neg f$ denotes the negation of $f$.
In addition to fixed-weight functions, we also attempt to study $\mathrm{D}_{\mathrm{ave}}(f)$ for circuits. For circuits consisting of AND, OR, NOT gates with unbounded fan-in, we show $\mathrm{D}_{\mathrm{ave}}(f)\leq 2s$, if $f$ can be computed by a circuit of $s$ gates. Note that $\mathrm{D}_{\mathrm{ave}}(\mathrm{XOR}_{n})=n$, and that $\mathrm{XOR}_{n}$ is computable by $O(n)$ gates; so the bound is almost tight.
Rossman introduced the notion of criticality, defined as the minimum value $\lambda\geq 1$ such that $\Pr_{\rho\sim\mathcal{R}_{p}}[\mathrm{D}(f|_{\rho})\geq t]<(p\lambda)^{t}$ for every $p\in[0,1]$ and $t\in\mathbb{N}$, where $\mathcal{R}_{p}$ denotes a $p$-random restriction, providing a framework to reformulate the switching lemmas. In terms of criticality, Håstad proved every width-$w$ CNF/DNF formula is $O(w)$-critical [Has86], and Rossman proved that every size-$s$ CNF/DNF formula is $O(\log s)$-critical [Ros17, Ros19]. Our next result is an attempt to prove the tightness of the switching lemmas [Has86, Ros17, Ros19], i.e., the corresponding upper bounds on criticality cannot be improved.
By applying an $(\frac{1}{2\lambda})$-random restriction as a query algorithm, we observe that $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{2\lambda})+O(1)$ for any $\lambda$-critical function $f$. As such, criticality upper bounds immediately imply upper bounds on $\mathrm{D}_{\mathrm{ave}}(f)$. For example, Håstad’s switching lemma implies that $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-1/O(w))$ for any CNF/DNF formula of width $w$ (Corollary 5.7); Rossman’s switching lemma implies that $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-1/O(\log s))$ for any CNF/DNF formula of size $s$ (Corollary 5.8).
On the other hand, if $\mathrm{D}_{\mathrm{ave}}(f)\geq n(1-1/\Omega(\lambda))$, then boolean function $f$ is $\Omega(\lambda)$-critical. In other words, lower bounds on $\mathrm{D}_{\mathrm{ave}}$ imply lower bounds on criticality. To address the tightness of the switching lemmas, we show there exist boolean functions with $\mathrm{D}_{\mathrm{ave}}(f)$ nearly matching the upper bounds (Corollary 5.7, 5.8), which are obtained straightforwardly from Håstad’s switching lemma [Has86] and Rossman’s switching lemma [Ros17, Ros19]. Our result implies that the criticality bounds $O(w)$ and $O(\log s)$ are tight up to a multiplicative $\log n$ factor.
Theorem 1.3.
For every integer $w\in[\log n+\log\log n+3,n]$, there exists some boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ computable by a DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$ such that
$\mathrm{D}_{\mathrm{ave}}(f)=n\left(1-\frac{\log n}{\Theta(w)}\right).$
Corollary 1.4.
For every integer $w\in[\log n+\log\log n+3,n]$, there exists an $\Omega(\frac{w}{\log n})$-critical DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$.
Besides the connection with criticality, it is meaningful to understand whether the upper bounds on $\mathrm{D}_{\mathrm{ave}}(f)$ (that follow from criticality upper bounds) are tight or not. For example, consider the bound $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{O(\log s)^{d-1}})$ for depth-$d$ size-$s$ $\mathsf{AC}^{0}$ circuits (Corollary 5.9). If the bound is tight, it would imply that random restriction is essentially the optimal average-case query algorithm for generic $\mathsf{AC}^{0}$ circuits. Otherwise, there exists a better query algorithm outperforming random restrictions. Either way, the answer would be interesting.
In Appendix B, we study (compute) the average-case query complexity (as well as some related average-case measures) for several functions, including the AND/OR/XOR functions, the penalty shoot-out function, symmetric functions, linear threshold functions and the tribes functions.
2 Preliminaries
Lemma 2.1 (Hoeffding’s inequality).
Let $X_{1},X_{2},\ldots,X_{m}$ be independent random variables such that $0\leq X_{i}\leq 1$, and let $S_{m}=X_{1}+X_{2}+\ldots+X_{m}$. For any $t>0$, we have
$\Pr[|S_{m}-\mathbb{E}[S_{m}]|\geq t]\leq 2\exp\left(-\frac{2t^{2}}{m}\right).$ (3)
The inequality (3) also holds when $X_{1},\ldots,X_{m}$ are obtained by sampling without replacement.
2.1 Boolean functions
A boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ maps each input $x\in\{0,1\}^{n}$ to 0 or 1. The weight, denoted by $\operatorname{wt}(f)$, is the number of inputs on which $f$ outputs 1. Let $\mathcal{B}_{n,m}=\{f:\{0,1\}^{n}\rightarrow\{0,1\}\mid\operatorname{wt}(f)=m\}$ denote the set of all $n$-variable boolean functions with weight $m$.
A restriction $\rho:\{1,\ldots,n\}\rightarrow\{0,1,\star\}$ is a mapping fixing some variables to 0 or 1. We write $f|_{\rho}$ for the subfunction of $f$ obtained by fixing its input by $\rho$. Let $\operatorname{supp}(\rho)=\rho^{-1}(\{0,1\})$ denote the support of the restriction $\rho$. The weight of $x$, denoted by $|x|$, is the number of 1’s in $x$. The bitwise negation of $x$ is denoted by $\neg x=(1-x_{1},\ldots,1-x_{n})$.
Let $F:\{0,1\}^{n}\rightarrow\{0,1\}$ and $G:\{0,1\}^{m}\rightarrow\{0,1\}$ be two boolean functions. Define the composition $F\circ G$ by
$(F\circ G)(x)=F(G(x^{(1)}),\ldots,G(x^{(n)}))$
for $x=(x^{(1)},\ldots,x^{(n)})\in\{0,1\}^{nm}$ and $x^{(i)}\in\{0,1\}^{m}$.
Define $x\preceq y$ if and only if $x_{i}\leq y_{i}$ for all $i\in\{1,2,\ldots,n\}$. Say $f$ is monotone if and only if $f(x)\leq f(y)$ for all inputs $x,y\in\{0,1\}^{n}$.
2.2 Decision trees
A (deterministic) decision tree $T$ is a binary tree. Each internal node is labeled by some integer $i\in\{1,2,\ldots,n\}$, and the edges and the leaves are labeled by 0 or 1. Repeatedly querying $x_{i}$ and following the edge labeled by $x_{i}$, the decision tree $T$ finally reaches a leaf and outputs the leaf’s label, called the value $T(x)$ of $T$ on input $x$. The cost of deciding the value $T(x)$, denoted by $\mathrm{cost}(T,x)$, is the length of the root-to-leaf path which $T$ passes through. The depth of $T$ is the maximum cost $\max_{x\in\{0,1\}^{n}}\mathrm{cost}(T,x)$. We say $T$ computes $f$ (with zero error) if $T(x)=f(x)$ for every $x\in\{0,1\}^{n}$. A query algorithm queries some variables and determines the value of the function; a query algorithm can be viewed as a family of decision trees.
A randomized decision tree $\mathcal{T}$ is a distribution over decision trees. Let $\operatorname{supp}(\mathcal{T})$ denote the support of $\mathcal{T}$. The cost on input $x\in\{0,1\}^{n}$, denoted by $\mathrm{cost}(\mathcal{T},x)$, is the expected number of variables queried by $\mathcal{T}$ on input $x$. The depth of $\mathcal{T}$ is the maximum cost $\max_{x\in\{0,1\}^{n}}\mathrm{cost}(\mathcal{T},x)$. We say $\mathcal{T}$ computes $f$ with zero error if every $T\in\operatorname{supp}(\mathcal{T})$ computes $f$. The randomized tree $\mathcal{T}$ computing $f$ is bounded-error if $T(x)\neq f(x)$ with probability no more than $1/3$ for every $x\in\{0,1\}^{n}$. (Note that the constant $1/3$ is not essential.)
2.3 Complexity measures
The (worst-case) deterministic query complexity of a boolean function $f$, denoted by $\mathrm{D}(f)$, is the minimum depth of decision trees computing $f$. The average-case deterministic query complexity of a boolean function $f$, denoted by $\mathrm{D}_{\mathrm{ave}}(f)$, is the minimum average depth of zero-error deterministic decision trees computing $f$ under a uniform input distribution.
Definition 2.2.
The average-case deterministic query complexity of $f:\{0,1\}^{n}\rightarrow\{0,1\}$ under a uniform distribution is defined by
$\mathrm{D}_{\mathrm{ave}}(f)=\min_{T}\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}[\mathrm{cost}(T,x)],$
where $T$ is taken over all zero-error deterministic decision trees computing $f$.
The following lemma says $\mathrm{D}_{\mathrm{ave}}(f)$ equals average-case zero-error randomized query complexity. Thus, we do not distinguish between these two.
Lemma 2.3 (Remark 8.63 in [ODo14]).
Define $\mathrm{ZR}_{\mathrm{ave}}(f)=\min_{\mathcal{T}}\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}[\mathrm{cost}(\mathcal{T},x)]$, where $\mathcal{T}$ is taken over all zero-error randomized decision trees computing $f$. We have $\mathrm{ZR}_{\mathrm{ave}}(f)=\mathrm{D}_{\mathrm{ave}}(f)$ for any boolean function $f$.
Proof.
Let $\mathcal{T}$ be an optimal zero-error randomized decision tree computing $f$ such that $\mathrm{ZR}_{\mathrm{ave}}(f)=\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}[\mathrm{cost}(\mathcal{T},x)]$. Since every deterministic decision tree can be viewed as a randomized tree, we have $\mathrm{ZR}_{\mathrm{ave}}(f)\leq\mathrm{D}_{\mathrm{ave}}(f)$.
On the other hand, there exists a deterministic tree $T$ drawn from $\mathcal{T}$ such that $\mathrm{D}_{\mathrm{ave}}(f)\leq\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}[\mathrm{cost}(T,x)]\leq\mathop{\mathbb{E}}_{T\sim\mathcal{R}}[\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}[\mathrm{cost}(T,x)]]=\mathrm{ZR}_{\mathrm{ave}}(f)$, and thus $\mathrm{D}_{\mathrm{ave}}(f)\leq\mathrm{ZR}_{\mathrm{ave}}(f)$. Therefore, we conclude $\mathrm{ZR}_{\mathrm{ave}}(f)=\mathrm{D}_{\mathrm{ave}}(f)$. ∎
The (worst-case) bounded-error randomized query complexity, denoted by $\mathrm{R}(f)$, is the minimum depth of all bounded-error randomized decision trees computing $f$. The average-case bounded-error randomized query complexity is defined by $\mathrm{R}_{\mathrm{ave}}(f)=\min_{\mathcal{T}}\mathop{\mathbb{E}}_{x\in\{0,1\}}[\mathrm{cost}(\mathcal{T},x)]$, where $\mathcal{T}$ is taken over all bounded-error randomized decision tree computing $f$ under the uniform input distribution.
Define $x^{\oplus i}=(x_{1},\ldots,x_{i-1},1-x_{i},x_{i+1},\ldots,x_{n})$. Say a variable $x_{i}$ is pivotal on input $x$ if $f(x)\neq f(x^{\oplus i})$. The sensitivity of function $f$ on input $x$, denoted by $\mathrm{sens}_{x}(f)$, is the number of pivotal variables on input $x$. The sensitivity of function $f$ is defined by $\mathrm{s}(f)=\max_{x\in\{0,1\}^{n}}\mathrm{sens}_{x}(f)$. The average sensitivity of function $f$ is defined by $\mathrm{s}_{\mathrm{ave}}(f)=\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}[\mathrm{sens}_{x}(f)]$, where $x$ is taken over the uniform distribution.
Let $x^{\oplus S}$ denote the vector by flipping the $i$-th bit for all $i\in S$. The block sensitivity of function $f$ on input $x$, denoted by $\mathrm{bs}_{x}(f)$, is the maximum integer $k$ such that there exist disjoint subsets $B_{1},\ldots,B_{k}\subseteq\{1,\ldots,n\}$ such that $f(x)\neq f(x^{\oplus B_{i}})$ for all $i=1,2,\ldots,k$. The block sensitivity of function $f$ is defined by $\mathrm{bs}(f)=\max_{x\in\{0,1\}^{n}}\mathrm{bs}_{x}(f)$. The average block sensitivity is defined by $\mathrm{bs}_{\mathrm{ave}}(f)=\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}[\mathrm{bs}_{x}(f)]$, where $x$ is over the uniform input distribution.
Given a boolean function $f:\{0,1\}^{n}\to\{0,1\}$, a certificate on input $x$ is a subset $S\subseteq\{1,\ldots,n\}$ such that $f(x)=f(y)$ for any input $y\in\{0,1\}^{n}$ satisfying $x_{i}=y_{i}$ for all $i\in S$. The certificate complexity $\mathrm{C}_{x}(f)$ on input $x$ is the size of the smallest certificate on input $x$. The certificate complexity of function $f$ is defined by $\mathrm{C}(f)=\max_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)$. The average certificate complexity is defined by $\mathrm{C}_{\mathrm{ave}}(f)=\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}[\mathrm{C}_{x}(f)]$, where $x$ is over a uniform input distribution.
Let $M_{1},M_{2}:\bigcup_{n\geq 0}\mathcal{B}_{n}\to\mathbb{R}_{\geq 0}$ be two functions (measures) on boolean functions, where $\mathcal{B}_{n}$ denote the set of all $n$-variable boolean functions. We say $M_{1}$ and $M_{2}$ are polynomially related if there exist constants $c_{1},c_{2}>0$ such that, when $n$ is sufficiently large, $M_{2}(f)\leq M_{1}(f)^{c_{1}}$ and $M_{1}(f)\leq M_{2}(f)^{c_{2}}$ for any $n$-variable boolean function $f$.
2.4 Circuits and formulas
A clause is a logical OR of variables or their negations, and a term is a logical AND of variables or their negations. A conjunctive normal form (CNF) formula is a logical AND of clauses, and a disjunctive normal form (DNF) formula is a logical OR of terms. The size of a CNF (respectively, DNF) formula is the number of the clauses (respectively, the terms). The width of a CNF (respectively, DNF) formula is the maximum variable number of the clauses (respectively, the terms).
A circuit $F$ is a directed acyclic graph with $n$ nodes of no incoming edge, called sources, and a node of no outgoing edge, called sink. Apart from the sources, the other nodes are called gates. Each gate is labeled by AND, OR or NOT, and each AND (respectively, OR, NOT) node computes the logical AND (respectively, OR, NOT) of the values of its incoming nodes. The fan-in of a gate is the number of its incoming edges, and the fan-out of a gate is the number of its outgoing edges. The fan-in of NOT gates is fixed to $1$. The size of $F$ is the number of the gates. The depth is the length of the longest path between the sink and the sources. Obviously, every CNF or DNF formula can be represented as a circuit. An $\mathsf{AC}^{0}$ circuit is a circuit of polynomial size, constant depth and AND/OR gates with unbounded fan-in. A formula is a circuit of gates with fan-out $1$.
3 Super-polynomial gaps
In query complexity, many worst-case complexity measures, including $\mathrm{D}(f)$, $\mathrm{R}(f)$, $\mathrm{C}(f)$, $\mathrm{bs}(f)$, $\mathrm{s}(f)$, etc., are known to be polynomially related (for example, see the summaries in [BD02, ABK16, ABK+21] and Huang’s proof of the Sensitivity Conjecture [Hua19]). In contrast, Ambainis and de Wolf demonstrated that $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$ and $\mathrm{Q}_{\mathrm{ave}}(f)$ are not polynomially related by showing exponential or super-exponential gaps [AD01]. In this section, we are going to show that, even for balanced monotone functions, polynomial relation does not hold between any two of the measures: $\mathrm{D}(f)$, $\mathrm{R}(f)$, $\mathrm{C}(f)$, $\mathrm{bs}(f)$, $\mathrm{s}(f)$, $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$, $\mathrm{C}_{\mathrm{ave}}(f)$, $\mathrm{bs}_{\mathrm{ave}}(f)$, $\mathrm{s}_{\mathrm{ave}}(f)$
Firstly, we have the following inequalities, most of which are straightforward by definition.
Proposition 3.1 ([AD01]).
$\mathrm{R}_{\mathrm{ave}}(f)\leq\mathrm{D}_{\mathrm{ave}}(f)\leq\mathrm{D}(f)$ for any boolean function $f$.
Proposition 3.2 ([Nis89, AD01]).
$\mathrm{R}_{\mathrm{ave}}(f)=\Omega(\mathrm{bs}_{\mathrm{ave}}(f))$ for any boolean function $f$.
Proposition 3.3.
$\mathrm{s}_{\mathrm{ave}}(f)\leq\mathrm{bs}_{\mathrm{ave}}(f)\leq\mathrm{C}_{\mathrm{ave}}(f)\leq\mathrm{D}_{\mathrm{ave}}(f)$ for any boolean function $f$.
Secondly, we separate the average-case measures $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$, $\mathrm{C}_{\mathrm{ave}}(f)$, $\mathrm{bs}_{\mathrm{ave}}(f)$, $\mathrm{s}_{\mathrm{ave}}(f)$ from the worst-case measures $\mathrm{D}(f)$, $\mathrm{R}(f)$, $\mathrm{C}(f)$, $\mathrm{bs}(f)$, $\mathrm{s}(f)$. It suffices to show a super-exponential gap between $\mathrm{D}(f)$ and $\mathrm{D}_{\mathrm{ave}}(f)$ because $\mathrm{D}_{\mathrm{ave}}(f)$ is the largest among these average-case measures by Proposition 3.1, 3.3, and all the worst-case measures are polynomially related.
Corollary 3.4 (Proposition B.4).
Let $n$ be odd, and let $f:\{0,1\}^{n}\to\{0,1\}$ be the penalty shoot-out function. We have (1) $f$ is monotone and balanced; (2) $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$; and (3) $\mathrm{D}(f)=n$.
Thirdly, there are monotone functions $f$ such that the gaps between any two of the average-case measures $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$, $\mathrm{C}_{\mathrm{ave}}(f)$, $\mathrm{bs}_{\mathrm{ave}}(f)$ and $\mathrm{s}_{\mathrm{ave}}(f)$ are super-polynomial. For example:
  • Let $f=\mathrm{Thr}_{n,\frac{n}{10}}$ be the threshold function. We have $\mathrm{D}_{\mathrm{ave}}(f)=\Omega(n)$ and $\mathrm{R}_{\mathrm{ave}}(f)=O(1)$ (Proposition B.7).
  • Let $f=\mathrm{Tribes}_{w,w2^{w}}$ be tribes function. We have $\mathrm{D}_{\mathrm{ave}}(f)=\Theta(2^{w})$, $\mathrm{R}_{\mathrm{ave}}(f)=\Theta(2^{w})$, $\mathrm{bs}_{\mathrm{ave}}(f)=\mathrm{C}_{\mathrm{ave}}(f)=\Theta(w)$ and $\mathrm{s}_{\mathrm{ave}}(f)=o(1)$ (Proposition B.19, B.20, B.21, B.24).
  • Let $f=\mathrm{Maj}_{n}\circ\mathrm{AND}_{3\log n}$. We have $\mathrm{C}_{\mathrm{ave}}(f)=\Omega(n)$ and $\mathrm{bs}_{\mathrm{ave}}(f)=O(1)$ (Lemma 3.10).
However, they are all biased. In order to exhibit the gaps among monotone balanced functions, we employ a “neutralization” trick to convert a biased monotone function into a balanced one.
Definition 3.5.
The dual of the function $f$ is defined by $f^{\dagger}(x)=1-f(\neg x)$.
The proof of the following proposition is straightforward by definition.
Proposition 3.6.
For any boolean function $f$, we have
  • (1)
    $f^{\dagger}$ is monotone if and only if $f$ is monotone;
  • (2)
    $\Pr_{x\in\{0,1\}^{n}}[f^{\dagger}(x)=1]=\Pr_{x\in\{0,1\}^{n}}[f(x)=0]$; and
  • (3)
    $\mathrm{D}_{\mathrm{ave}}(f)=\mathrm{D}_{\mathrm{ave}}(f^{\dagger})$, $\mathrm{R}_{\mathrm{ave}}(f)=\mathrm{R}_{\mathrm{ave}}(f^{\dagger})$, $\mathrm{C}_{\mathrm{ave}}(f)=\mathrm{C}_{\mathrm{ave}}(f^{\dagger})$, $\mathrm{bs}_{\mathrm{ave}}(f)=\mathrm{bs}_{\mathrm{ave}}(f^{\dagger})$, $\mathrm{s}_{\mathrm{ave}}(f)=\mathrm{s}_{\mathrm{ave}}(f^{\dagger})$.
Lemma 3.7.
Suppose $f$ is monotone. Let $g(x,y,z)=\mathrm{Maj}_{3}(f(x),f^{\dagger}(y),z)$ for $x,y\in\{0,1\}^{n}$ and $z\in\{0,1\}$. Then, $g$ is a monotone balanced function such that
  • $\mathrm{D}_{\mathrm{ave}}(g)=\Theta(\mathrm{D}_{\mathrm{ave}}(f))$,
  • $\mathrm{R}_{\mathrm{ave}}(g)=\Theta(\mathrm{R}_{\mathrm{ave}}(f))$,
  • $\mathrm{C}_{\mathrm{ave}}(g)=\Theta(\mathrm{C}_{\mathrm{ave}}(f))$,
  • $\mathrm{bs}_{\mathrm{ave}}(g)=\Theta(\mathrm{bs}_{\mathrm{ave}}(f))$, and
  • $\mathrm{s}_{\mathrm{ave}}(g)=\max\{\Theta(\mathrm{s}_{\mathrm{ave}}(f)),\Theta(1)\}$.
Proof.
The function $g$ is balanced, since
$\Pr_{x,y,z}[g(x,y,z)=1]=$ $\ \Pr_{x,y,z}[(f(x),f^{\dagger}(y),z)=011]+\Pr_{x,y,z}[(f(x),f^{\dagger}(y),z)=101]$
$\ +\Pr_{x,y,z}[(f(x),f^{\dagger}(y),z)=110]+\Pr_{x,y,z}[(f(x),f^{\dagger}(y),z)=111]$
$=$ $\ \frac{1}{2}(1-p)^{2}+\frac{1}{2}p^{2}+\frac{1}{2}p(1-p)+\frac{1}{2}p(1-p)=\frac{1}{2},$
where $p=\Pr_{x\in\{0,1\}^{n}}[f(x)=1]$. The function $g$ is monotone, since $\mathrm{Maj}_{3},f,f^{\dagger},z$ are all monotone. The value $g(x,y,z)$ can be determined by applying the optimal algorithms for $f(x),f^{\dagger}(y)$, and querying the variable $z$. Thus, $\mathrm{D}_{\mathrm{ave}}(g)\leq\mathrm{D}_{\mathrm{ave}}(f)+\mathrm{D}_{\mathrm{ave}}(f^{\dagger})+1=O(\mathrm{D}_{\mathrm{ave}}(f))$. Notice that
$g(x,y,z)=\begin{cases}0,&\quad f^{\dagger}(y)=0\wedge z=0\\ 1,&\quad f^{\dagger}(y)=1\wedge z=1\\ f(x),&\quad\text{otherwise}\\ \end{cases}.$
So, $g(x,y,z)=f(x)$ if $f^{\dagger}(y)\neq z$. Ignoring the costs of querying $y$ and $z$, we have
$\mathrm{D}_{\mathrm{ave}}(g)\geq\Pr_{y,z}[f^{\dagger}(y)\neq z]\cdot\mathrm{D}_{\mathrm{ave}}(f)=\frac{1}{2}\mathrm{D}_{\mathrm{ave}}(f).$
We conclude $\mathrm{D}_{\mathrm{ave}}(g)=\Theta(\mathrm{D}_{\mathrm{ave}}(f))$. A similar argument works for $\mathrm{R}_{\mathrm{ave}}$ and $\mathrm{C}_{\mathrm{ave}}$.
Similarly, since $g(x,y,z)=f(x)$ with probability at least $1/2$, we have $\mathrm{bs}_{\mathrm{ave}}(g)=\Omega(\mathrm{bs}_{\mathrm{ave}}(f))$. On the other hand, we shall prove $\mathrm{bs}_{u}(g)\leq\mathrm{bs}_{x}(f)+\mathrm{bs}_{y}(f^{\dagger})+1$ for every input $u=(x,y,z)\in\{0,1\}^{2n+1}$, which implies $\mathrm{bs}_{\mathrm{ave}}(g)=O(\mathrm{bs}_{\mathrm{ave}}(f))$. Let $m=\mathrm{bs}_{u}(g)$. So there exist $m$ disjoint subsets $U_{1},\ldots,U_{m}\subseteq\{1,\ldots,2n+1\}$ such that $g(u)\neq g(u^{\oplus U_{i}})$ for all $i=1,\ldots,m$.
Assume the input variable $z$ belongs to $U_{m}$ without loss of generality. For any $i\leq m-1$, flipping the bits in each $U_{i}$ should be able to alter $f(x)$ or $f^{\dagger}(y)$, since $g(u)=\mathrm{Maj}_{3}(f(x),f^{\dagger}(y),z)$ and $z$ does not belong to $U_{1},\ldots,U_{m}$. In other words, we have $f(x)\neq f(x^{\oplus X_{i}})$ or $f^{\dagger}(y)\neq f^{\dagger}(y^{\oplus Y_{i}})$, where $X_{i}=U_{i}\cap\{1,\ldots,n\}$ and $Y_{i}=U_{i}\cap\{n+1,\ldots,2n\}$, for all $i=1,\ldots,m-1$. Note that $X_{i}$ is pivotal on input $x$ for $f$ if $f(x)\neq f(x^{\oplus X_{i}})$, and that $Y_{i}$ is pivotal on input $y$ for $f^{\dagger}$ if $f^{\dagger}(y)\neq f^{\dagger}(y^{\oplus Y_{i}})$.
If $m>\mathrm{bs}_{x}(f)+\mathrm{bs}_{y}(f^{\dagger})+1$, then we can find $\mathrm{bs}_{x}(f)+1$ pivotal blocks for $f$ on input $x$ or $\mathrm{bs}_{y}(f^{\dagger})+1$ pivotal blocks for $f^{\dagger}$ on input $y$, contradicting the maximality of $\mathrm{bs}_{x}(f)$ or $\mathrm{bs}_{y}(f^{\dagger})$. Thus, we conclude that $m=\mathrm{bs}_{u}(g)\leq\mathrm{bs}_{x}(f)+\mathrm{bs}_{y}(f^{\dagger})+1$ for all $u=(x,y,z)\in\{0,1\}^{2n+1}$.
For any $i=1,\ldots,n$, we have $g(u)\neq g(u^{\oplus i})$ if and only if $f(x)\neq f(x^{\oplus i})$ and $f^{\dagger}(y)\neq z$, since $g(u)$ will not change if $f^{\dagger}(y)=z$. Similarly, for $i=1,\ldots,n$, we have $g(u)\neq g(u^{\oplus i^{*}})$ if and only if $f(x)\neq z$ and $f^{\dagger}(y)\neq f^{\dagger}(y^{\oplus i})$, where $i^{*}=i+n\in\{n+1,\ldots,2n\}$. Therefore,
$\mathrm{s}_{\mathrm{ave}}(g)=$ $\ \sum_{i=1}^{2n+1}\Pr_{u\in\{0,1\}^{2n+1}}\left[g(u)\neq g(u^{\oplus i})\right]$
$=$ $\ \sum_{i=1}^{n}\left(\Pr_{x,y,z}\left[\left(f(x)\neq f(x^{\oplus i})\right)\wedge\left(f^{\dagger}(y)\neq z\right)\right]+\Pr_{x,y,z}\left[\left(f(x)\neq z\right)\wedge\left(f^{\dagger}(y)\neq f^{\dagger}(y^{\oplus i})\right)\right]\right)$
$\ +\Pr_{x,y,z}\left[g(x,y,z)\neq g(x,y,z\oplus 1)\right]$
$=$ $\ \Pr_{y,z}[f^{\dagger}(y)\neq z]\cdot\mathrm{s}_{\mathrm{ave}}(f)+\Pr_{x,z}[f(x)\neq z]\cdot\mathrm{s}_{\mathrm{ave}}(f^{\dagger})+\Pr_{x,y,z}\left[g(x,y,z)\neq g(x,y,z\oplus 1)\right].$ (4)
The equation (4) implies that $\mathrm{s}_{\mathrm{ave}}(g)=\Theta(\mathrm{s}_{\mathrm{ave}}(f))$ if $\mathrm{s}_{\mathrm{ave}}(f)=\Omega(1)$, and that $\mathrm{s}_{\mathrm{ave}}(f)\leq 1+o(1)=O(1)$ if $\mathrm{s}_{\mathrm{ave}}(f)=o(1)$. Since $\Pr_{x,y,z}\left[g(x,y,z)\neq g(x,y,z\oplus 1)\right]=\Pr_{x,y}[f(x)\neq f^{\dagger}(y)]=p^{2}+(1-p)^{2}\geq 1/2$, the average sensitivity $\mathrm{s}_{\mathrm{ave}}(g)=\Omega(1)$. Thus, $\mathrm{s}_{\mathrm{ave}}(g)=\max\{\Theta(\mathrm{s}_{\mathrm{ave}}(f)),\Theta(1)\}$. ∎
Finally, we are ready to show super-polynomial gaps for monotone balanced functions. In fact, by Proposition 3.1, 3.3, it suffices to show 5 gaps between (1) $\mathrm{s}_{\mathrm{ave}}(f)$ and $\mathrm{bs}_{\mathrm{ave}}(f)$; (2) $\mathrm{bs}_{\mathrm{ave}}(f)$ and $\mathrm{C}_{\mathrm{ave}}(f)$; (3) $\mathrm{C}_{\mathrm{ave}}(f)$ and $\mathrm{R}_{\mathrm{ave}}(f)$; (4) $\mathrm{C}_{\mathrm{ave}}(f)$ and $\mathrm{D}_{\mathrm{ave}}(f)$; and (5) $\mathrm{R}_{\mathrm{ave}}(f)$ and $\mathrm{D}_{\mathrm{ave}}(f)$.
Corollary 3.8.
Let $f=\mathrm{Thr}_{n,\frac{n}{10}}$ and $g(x,y,z)=\mathrm{Maj}_{3}(f(x),f^{\dagger}(y),z)$ for $x,y\in\{0,1\}^{n}$ and $z\in\{0,1\}$. Then, $\mathrm{R}_{\mathrm{ave}}(g)=O(1)$ and $\mathrm{D}_{\mathrm{ave}}(g)=\Omega(n)$
Proof.
The result follows from Lemma 3.7 and Proposition B.7. ∎
Corollary 3.9.
Let $s=w2^{w}$ and $f=\mathrm{Tribes}_{w,s}$ The function $g(x,y,z)=\mathrm{Maj}_{3}(f(x),f^{\dagger}(y),z)$ has $\mathrm{s}_{\mathrm{ave}}(g)=\Theta(1)$ and $\mathrm{bs}_{\mathrm{ave}}(g)=\mathrm{C}_{\mathrm{ave}}(g)=\Theta(w)$ and $\mathrm{R}_{\mathrm{ave}}(g)=\Theta(2^{w})$ and $\mathrm{D}_{\mathrm{ave}}(g)=\Theta(2^{w})$ for $x,y\in\{0,1\}^{ws}$ and $z\in\{0,1\}$.
Proof.
By Proposition B.19, we have $\mathrm{D}_{\mathrm{ave}}(f)=2(2^{w}-1)(1-(1-\frac{1}{2^{w}})^{w2^{w}})=\Theta(2^{w})$. Then, we have $\mathrm{R}_{\mathrm{ave}}(f)=\Theta(2^{w})$ by Proposition B.24, since $s\gg 2^{w}$. Moreover, we have $\mathrm{s}_{\mathrm{ave}}(f)=\frac{2w^{2}2^{w}}{2^{w}-1}(1-\frac{1}{2^{w}})^{w2^{w}}=O(\frac{w^{2}}{e^{w}})=o(1)$ by Proposition B.20, and $\mathrm{bs}_{\mathrm{ave}}(f)=\mathrm{C}_{\mathrm{ave}}(f)=w(1-(1-\frac{1}{2^{w}})^{w2^{w}})+w2^{w}(1-\frac{1}{2^{w}})^{w2^{w}}=\Theta(w)$ by Proposition B.21. Finally, the corollary follows from Lemma 3.7. ∎
Lemma 3.10.
Let $f=\mathrm{Maj}_{n}\circ\mathrm{AND}_{3\log n}$. We have $\mathrm{bs}_{\mathrm{ave}}(f)=O(1)$ and $\mathrm{C}_{\mathrm{ave}}(f)=\Omega(n)$.
Proof.
Let $t=3\log n$. Since every certificate contains either $\lceil\frac{n+1}{2}\rceil$ zeros or $\lceil\frac{n+1}{2}\rceil$ ones at least, we have $\mathrm{C}_{x}(f)=\Omega(n)$ for all inputs $x\in\{0,1\}^{nt}$. Thus, $\mathrm{C}_{\mathrm{ave}}(f)=\Omega(n)$.
Rewrite $f(x)=\mathrm{Maj}_{n}(u(x))$ for $u(x)=(u_{1}(x),\ldots,u_{n}(x))$ and $u_{i}(x)=\bigwedge_{j=1}^{t}x_{i,j}$. Let $0^{n}$ denote the all-zero vector of length $n$. The $\mathrm{bs}_{0^{n}}(\mathrm{Maj}_{n})$ equals $1$, since we have to flip more than $n/2$ bits to alter the value $\mathrm{Maj}_{n}(0^{n})$. Since $\Pr_{x\in\{0,1\}^{nt}}[u(x)\neq 0^{n}]=1-(1-\frac{1}{2^{(3\log n)}})^{n}\leq n^{-2}$, we have
$\mathrm{bs}_{\mathrm{ave}}(\mathrm{Maj}_{n}\circ\mathrm{AND}_{t})\leq\Pr_{x\in\{0,1\}^{nt}}[u(x)=0^{n}]\cdot 1+\Pr_{x\in\{0,1\}^{nt}}[u(x)\neq 0^{n}]\cdot(3n\log n)\leq 1+\frac{3\log n}{n}.$
Corollary 3.11.
Let $f=\mathrm{Maj}_{n}\circ\mathrm{AND}_{3\log n}$ and $g(x,y,z)=\mathrm{Maj}_{3}(f(x),f^{\dagger}(y),z)$ for $x,y\in\{0,1\}^{3n\log n}$ and $z\in\{0,1\}$. Then, $g$ is a balanced monotone function such that $\mathrm{bs}_{\mathrm{ave}}(g)=O(1)$ and $\mathrm{C}_{\mathrm{ave}}(g)=\Omega(n)$.
Proof.
The corollary follows from Lemma 3.7, 3.10. ∎
4 Fixed-weight functions
4.1 Upper bound
As a warm-up, we show a simple upper bound $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\operatorname{wt}(f)+O(1)$ for any boolean function $f$. Say a query algorithm is reasonable if it stops querying as long as the subfunction is constant. For convenience, we say input $x$ is a black point if $f(x)=1$.
Lemma 4.1.
$\mathrm{D}_{\mathrm{ave}}(f)\leq\log\operatorname{wt}(f)+2$ for any non-constant boolean function $f$.
Proof.
Let $m=\operatorname{wt}(f)$. We prove Lemma 4.1 by induction on $m$. When $m=1$, we have $\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})\leq 2$ by Proposition B.1. Assume the variable $x_{i}$ is queried first. Let $m_{0}=\operatorname{wt}(f|_{x_{i}=0})$ and $m_{1}=\operatorname{wt}(f|_{x_{i}=1})$. If $m_{b}=0$, where $b\in\{0,1\}$, any reasonable algorithm will stop on a constant subfunction $f|_{x_{i}=b}$. So the $\mathrm{D}_{\mathrm{ave}}(f)\leq 1+\frac{1}{2}(\log m+2)\leq\log m+2$ by the induction hypothesis. Otherwise, by the induction hypothesis and the AM-GM inequality, we have
$\mathrm{D}_{\mathrm{ave}}(f)$ $\leq 1+\frac{1}{2}(\log m_{0}+2)+\frac{1}{2}(\log m_{1}+2)$
$=\log(2\sqrt{m_{0}m_{1}})+2$
$\leq\log m+2.$
To prove an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$, it suffices to design a query algorithm computing $f$. The query algorithm we design for proving Lemma 4.1 and Theorem 1.1 are based on the same approach: query an arbitrary variable until the subfunction meets certain border conditions. The border conditions in the proofs:
  • Lemma 4.1: $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$ if $\operatorname{wt}(f)=1$ (Proposition B.1).
  • Theorem 1.1: $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$ if $\operatorname{wt}(f)\leq 4\log n$ (Corollary 4.5).
To prove the second one, let us fix the $\operatorname{wt}(f)$ and put the black points in a $\operatorname{wt}(f)\times n$ table with each row representing a black point. Notice the fixed number of distinct columns up to $2^{\operatorname{wt}(f)}$. When $n\gg 2^{\operatorname{wt}(f)}$, numerous duplicate columns are identified by the pigeonhole principle. Therefore, numerous black points share the same bit on multiple coordinates, leading to the construction of a decision tree with $O(1)$ average depth.
Suppose $f\in\mathcal{B}_{n,m}$ has $m$ black points $x^{(1)},\ldots,x^{(m)}\in\{0,1\}^{n}$. We call $c_{i}=(x_{i}^{(1)},\ldots,x_{i}^{(m)})$ the black column of the coordinate $i$. Say two coordinates $i,j$ are positively correlated if $c_{i}=c_{j}$ or negatively correlated if $c_{i}=\neg c_{j}$. An equivalent coordinate set (ECS) with respect to $f$ is a set of correlated coordinates. Say a coordinate set $S\subseteq\{1,\ldots,n\}$ is pure if $c_{i}$ is an all-zero or all-one vector for each $i\in S$; otherwise, say $S$ is mixed. For example, in Table 1, the set $\{5,9,{11}\}$ is pure, since $c_{5}=c_{9}=(0,0)$ and $c_{11}=(1,1)$, and the set $\{1,2,3\}$ is mixed, since $c_{1}=c_{3}=(0,1)$ and $c_{2}=(1,0)$.
black points $x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ $x_{5}$ $x_{6}$ $x_{7}$ $x_{8}$ $x_{9}$ $x_{10}$ $x_{11}$
$x^{(1)}$ 0 1 0 1 0 1 0 0 0 0 1
$x^{(2)}$ 1 0 1 0 0 1 0 1 0 1 1
Table 1: A 11-variable boolean function with weight $2$.
Proposition 4.2.
Fix a boolean function $f$. Suppose $i,j$ are positively correlated (respectively, negatively correlated). For every input $x\in\{0,1\}^{n}$, if $f(x)=1$, then $x_{i}=x_{j}$ (respectively, $x_{i}\neq x_{j}$).
Proposition 4.3.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function, and $m=\operatorname{wt}(f)$. If $n>k\cdot 2^{m-1}$, there exists some maximal ECS of size $k+1$ at least.
Proposition 4.3 follows from the pigeonhole principle, since there are $2^{m-1}$ distinct maximal ECSs. Combining these facts, we can construct a “shallow” decision tree computing $f$ when $n$ is exponentially large.
Lemma 4.4.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function. If $n>2^{\operatorname{wt}(f)}$, then $\mathrm{D}_{\mathrm{ave}}(f)\leq 5$.
Proof.
Let $m=\operatorname{wt}(f)$. We prove $\mathrm{D}_{\mathrm{ave}}(f)\leq 5$ by induction on $n$. By Proposition 4.3, there exists a maximal ECS $I=\{i_{1},\ldots,i_{k}\}$ for some $k\geq 3$. Assume $i_{1},\ldots,i_{k}$ are positively correlated without loss of generality. For any black point $x$, by Proposition 4.2, we have $x_{i_{1}}=\cdots=x_{i_{k}}$.
If $|I|=n$, any $f$’s black point $x$ satisfies $x_{1}=\cdots=x_{n}$. So there are at most two black points, all-zero or all-one. By Lemma 4.1, $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\operatorname{wt}(f)+2\leq 3$. Thus, we can assume $|I|<n$, and thus there exists a coordinate $j\not\in I$.
Case 1: The ECS $I$ is mixed. Without loss of generality, assume $c_{i_{1}}\in\{0,1\}^{m}$ is not all-zero or all-one. Our query algorithm $T_{1}$ runs as follows (see Figure 1(a)):
  • (1)
    query $x_{i_{1}},x_{i_{2}}$;
  • (2)
    output 00 if $x_{i_{1}}\neq x_{i_{2}}$;
  • (3)
    if $x_{i_{1}}=x_{i_{2}}$, query $x_{j}$, and apply an optimal query algorithm to the subfunction.
The decision tree $T_{1}$ computes $f$, since input $x$ cannot be a black point if $x_{i_{1}}\neq x_{i_{2}}$. Note that the probability of querying exactly 2 variables is $\frac{1}{2}$, and the probability of querying at least 3 variables is $\frac{1}{2}$.
(a) Case 1: The ECS $I$ is mixed and $|I|<n$
(b) Case 2: The ECS $I$ is pure and $|I|<n$
Figure 1: The decision trees with $O(1)$ average depth.
For convenience, denote by $\rho_{u,v}$ the restriction that fixes $x_{i_{1}},x_{i_{2}}$ to $u\in\{0,1\}$, fixes $x_{r}$ to $v\in\{0,1\}$, and leaves other variables undetermined. Since $c_{i_{1}}$ is neither all-zero nor all-one, we cut out at least one black point after querying $x_{i_{1}}$. For example, consider the function $g$ given in Table 1. After querying $x_{1}$, we split Table 1 into two subfunctions $g|_{x_{1}=0}$ and $g|_{x_{1}=1}$, one with $x^{(1)}$ and the other with $x^{(2)}$, since $c_{1}=(0,1)$ contains both 00 and $1$. Similarly, after querying $x_{i_{1}}$, we split the table of $f$’s black points into two non-empty tables, since $c_{i_{1}}$ is neither all-zero nor all-one. Thus, we claim $\operatorname{wt}(f_{\rho_{u,v}})\leq m-1$ for any $u,v\in\{0,1\}$.
For any $u,v\in\{0,1\}$, the number of the input variables of $f|_{\rho_{u,v}}$ is $n-3>2^{m}-3>2^{m-1}$ for any $m\geq 3$. By the induction hypothesis, we have $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{u,v}})\leq 5$. Thus, $T_{1}$’s average cost is at most
$\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}\left[\mathrm{cost}(T_{1},x)\right]\leq\frac{1}{2}\cdot 2+\Pr_{x}\left[\mathrm{cost}(T_{1},x)\geq 3\right]\cdot(3+5)=5.$
Case 2: The ECS $I$ is pure. Assume $c_{i_{1}},\ldots,c_{i_{k}}\in\{0,1\}^{m}$ are all-zero vectors without loss of generality. By Proposition 4.2, any black point $x$ must satisfy $x_{i_{1}}=\cdots=x_{i_{k}}=0$. Our query algorithm $T_{2}$ runs as follows (see Figure 1(b)):
  • (1)
    query $x_{i_{1}}$, and output 00 if $x_{i_{1}}=1$;
  • (2)
    query $x_{i_{2}}$, and output 00 if $x_{i_{2}}=1$;
  • (3)
    query $x_{j}$, and apply an optimal query algorithm to the subfunction $f|_{\rho_{0,0}}$ or $f|_{\rho_{0,1}}$.
Note that the probability of querying exactly $b\leq 2$ variables is $2^{-b}$, and that the probability of querying at least 3 variables is $1/4$.
Since $I$ is pure and maximal, the coordinate $j\not\in I$ belongs to some mixed ECS, and $c_{j}$ is neither all-zero nor all-one. Then, we have $\operatorname{wt}(f|_{\rho_{0,0}})\leq m-1$ and $\operatorname{wt}(f|_{\rho_{0,1}})\leq m-1$. Since the variable number of $\operatorname{wt}(f|_{\rho_{0,v}})$ is $n-3>2^{m}-3\geq 2^{m-1}$ for any $m\geq 3$, by the induction hypothesis, $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{0,v}})\leq 5$ for any $v\in\{0,1\}$. Thus, $T_{2}$’s average cost is at most
$\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}\left[\mathrm{cost}(T,x)\right]\leq\frac{1}{2}\cdot 1+\frac{1}{4}\cdot 2+\Pr_{x}[\mathrm{cost}(T_{2},x)\geq 3]\cdot(3+5)=3.$
Summarizing the two cases, we conclude that $\mathrm{D}_{\mathrm{ave}}(f)\leq\max\{5,3\}=5$. ∎
Corollary 4.5.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function. If $\operatorname{wt}(f)\leq 4\log n$, then $\mathrm{D}_{\mathrm{ave}}(f)\leq 40$.
Proof.
Observe that $\mathrm{D}_{\mathrm{ave}}(f_{1}(x)\vee\cdots\vee f_{k}(x))\leq\mathrm{D}_{\mathrm{ave}}(f_{1})+\cdots+\mathrm{D}_{\mathrm{ave}}(f_{k})$ for any $f_{1},f_{2},\ldots,f_{k}$. Because one can query $f_{1}(x)$, $f_{2}(x)$, $\ldots$, $f_{k}(x)$ one by one, and compute $f_{1}(x)\vee\cdots\vee f_{k}(x)$ at the end; the expected number of variables queried is at most the sum of the expectations.
Let $k=8$. Divide the black points of $f$ evenly into $k$ disjoint parts $f_{1},\ldots,f_{k}$, each with at most $\lceil\frac{4\log n}{8}\rceil<\frac{4\log n}{8}+1$ black points. In other words, $f(x)=f_{1}(x)\vee\cdots\vee f_{k}(x)$, and $\operatorname{wt}(f_{i})\leq\lceil\frac{4\log n}{8}\rceil$.
Note that $\operatorname{wt}(f_{i})\leq\lceil\frac{4\log n}{8}\rceil\leq\frac{4\log n}{8}+1\leq\log n$ when $n\geq 4$. (When $n<4$, the lemma obviously holds.) Thus, by Lemma 4.4, $\mathrm{D}_{\mathrm{ave}}(f_{i})\leq 5$ for any $i$. Therefore, $\mathrm{D}_{\mathrm{ave}}(f)\leq\sum_{i=1}^{8}\mathrm{D}_{\mathrm{ave}}(f_{i})\leq 40$. ∎
Now, we are ready to prove Theorem 1.1.
Proof of Theorem 1.1.
Let $m=\operatorname{wt}(f)$. If $m\leq 4\log n$, we have $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$ by Corollary 4.5. Assume $m>4\log n$. Using induction on $n$, we shall prove
$\mathrm{D}_{\mathrm{ave}}(f)\leq\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+C,$ (5)
for any $n$-variable boolean function with weight $m\geq 4\cdot\log n$, where $C>0$ is a sufficiently large constant.
First, let us prove, if $m\geq n$, Lemma 4.1 immediately implies (5). Observe that $\log m+2\leq\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+C$ holds if and only if $m\geq n^{2^{-(C-2)}}\log n$. When $C\geq 4$, it is obvious that $m\geq n^{2^{-(C-2)}}\log n$. Thus, (5) follows from Lemma 4.1.
Our query algorithm is as follows:
  • If $\operatorname{wt}(f)\geq n$, we apply Lemma 4.1, which implies (5) as we have shown.
  • If $\operatorname{wt}(f)\leq 4\log n$, we apply the query algorithm in Corollary 4.5, which implies $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$.
  • Otherwise, query arbitrary $\ell=\lceil\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+3\rceil<\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+4$ variables, say, $x_{1},\ldots,x_{\ell}$.
  • Given the value of $x_{1},x_{2},\ldots,x_{\ell}$, say, $x_{1}=c_{1},\ldots,x_{\ell}=c_{\ell}$, apply our algorithm recursively on the subfunction $f|_{x_{1}=c_{1},\ldots,x_{\ell}=c_{\ell}}$.
Let $\rho_{1},\ldots,\rho_{2^{\ell}}$ enumerate all restrictions that fix $x_{1},\ldots,x_{\ell}$, while leaving all other variables undetermined. Let $\mathcal{R}$ be the uniform distribution over $\rho_{1},\ldots,\rho_{2^{\ell}}$. Note that $\operatorname{wt}(f)=\sum_{i=1}^{2^{\ell}}\operatorname{wt}(f|_{\rho_{i}})$, and that $\mathop{\mathbb{E}}_{\rho\sim\mathcal{R}}[\operatorname{wt}(f|_{\rho})]=\frac{m}{2^{\ell}}\geq\frac{\log n}{16\log\frac{m}{\log n}}$. By Markov’s inequality,
$\Pr_{\rho\sim\mathcal{R}}\left[\operatorname{wt}(f|_{\rho})>4\log\frac{m}{\log n}\mathop{\mathbb{E}}_{\rho\sim\mathcal{R}}[\operatorname{wt}(f|_{\rho})]\geq\frac{1}{4}\log n\right]\leq\frac{1}{4\log\frac{m}{\log n}}.$
According to the weight of $f|_{\rho_{i}}$, let us upper bound $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})$ case by case.
Case 1: $\operatorname{wt}(f|_{\rho_{i}})\leq 4\log n$. Since $m\leq n$, we have $\ell\leq\log n+\log\log n+4$. So the number of variables in $f|_{\rho_{i}}$ is $n-\ell\geq n-(\log n+\log\log n+4)>\sqrt{n}$ (when $n$ is large enough). Notice that $\operatorname{wt}(f|_{\rho_{i}})\leq O(\log(n-\ell))$. Thus, by Corollary 4.5, we have $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})\leq C_{1}$, where $C_{1}>0$ is an absolute constant.
Case 2: $\operatorname{wt}(f|_{\rho_{i}})>4\log n$. Note that $\operatorname{wt}(f|_{\rho_{i}})\geq 4\log n>4\log(n-\ell)$. By the induction hypothesis, we have
$\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})$ $\leq\log\frac{m}{\log(n-\ell)}+\log\log\frac{m}{\log(n-\ell)}+C$
$\leq 2\log\frac{m}{\log(n-\ell)}+C$
$\leq 2\log\frac{m}{\log n}+2+C,$
where the second step is because $\log\log\frac{m}{\log(n-\ell)}\leq\log\frac{m}{\log(n-\ell)}$, and the third step is because $n-\ell\geq\sqrt{n}$.
Averaging over all $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})$, we have
$\ \mathrm{D}_{\mathrm{ave}}(f)$
$\leq$ $\ \ell+\mathop{\mathbb{E}}_{\rho\sim\mathcal{R}}[\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})]$
$\leq$ $\ \ell+\Pr_{\rho_{i}\sim\mathcal{R}}\left[\operatorname{wt}(f|_{\rho})\leq\frac{1}{4}\log n\right]\cdot 40+\Pr_{\rho\sim\mathcal{R}}\left[\operatorname{wt}(f|_{\rho})>\frac{1}{4}\log n\right]\cdot{\left(2\log\frac{m}{\log n}+2+C\right)}$
$\leq$ $\ \log\frac{m}{\log n}+\log\log\frac{m}{\log n}+4+40+\frac{1}{4\log\frac{m}{\log n}}\cdot{\left(2\log\frac{m}{\log n}+2+C\right)}$
$\leq$ $\ \log\frac{m}{\log n}+\log\log\frac{m}{\log n}+45+\frac{C}{4}$
$\leq$ $\ \log\frac{m}{\log n}+\log\log\frac{m}{\log n}+C,$
if $C\geq 60$.
If $m\geq 4\log n$, the right hand side of (5) is at most $\log\frac{m}{\log n}+O(\log\log\frac{m}{\log n})$, completing the proof of Theorem 1.1. ∎
4.2 Lower bound
In this section, we are going to prove Theorem 1.2: $\mathrm{D}_{\mathrm{ave}}(f)\geq\log\frac{m}{\log n}-O(\log\log\frac{m}{\log n})$ for almost all functions $f\in\mathcal{B}_{n,m}$, showing Theorem 1.1 is tight up to an additive logarithmic term.
To illustrate the idea of the proof for Theorem 1.2, let us take the $\mathrm{XOR}_{n}$ function as an example. No matter which variable is queried next, the black points will be evenly partitioned, and the weight of the subfunction is exactly the half. Since the weight of the $\mathrm{XOR}_{n}$ function is $2^{n-1}$ and the algorithm cannot stop until the subfunction becomes constant, it has to query all $n$ variables on every input, which implies that $\mathrm{D}_{\mathrm{ave}}(\mathrm{XOR}_{n})=n$. Similarly, the main idea of our proof is to show almost all boolean functions have a similar property: regardless of which variable is queried next, all the black points are divided into two halves of nearly equal size. In other words, for almost all $f\in\mathcal{B}_{n,m}$, we shall prove the weight $\operatorname{wt}(f|_{P})$ is close to $\operatorname{wt}(f)\cdot 2^{-k}$ for any tree path $P$ that queries $k=\epsilon\log m$ variables.
[…]
  1. (1)
    Let $y^{(1)},\ldots,y^{(m)}\in\{0,1,\star\}^{n}$ be $m$ vectors, where, initially, all the variables are undetermined and marked by $\star$.
  2. (2)
    First, we draw $t_{0}=m$ balls, one by one, from a box with $2^{n-1}$ 00’s and $2^{n-1}$ $1$’s by sampling without replacement. Place these balls in the $i_{1}$-th position of each vector $y^{(1)},\ldots,y^{(m)}$ in order. Then, we discard the vectors in which ($(\neg v_{1})$-ball is placed. Suppose that there are $t_{1}\leq t_{0}$ vectors remaining. (Why $2^{n-2}$ instead of $2^{n-1}$? Because $\neg v_{1}$ balls have been discarded.)
  3. (3)
    Next, we draw $t_{1}$ balls from a box with $2^{n-2}$ zeros and $2^{n-2}$ ones, put them in the $i_{2}$-th position of the remaining vectors, and discard the vectors in which ($(\neg v_{2})$-ball is placed.
  4. (4)
    Repeat steps 2) and 3) for $k$ steps in total.
  5. (5)
    Output the number of the remaining vectors in the $k$-th step.
For sampling $f\in\mathcal{B}_{n,m}$ uniformly at random, a straightforward approach proceeds as follows: (1) select $m$ inputs $x^{(1)},\cdots,x^{(m)}\in\{0,1\}^{n}$ uniformly at random without replacement; (2) set $f(x^{(i)})=1$ for all $i=1,2,\ldots,m$; and (3) set the value of the other inputs to 00. This can also be done by repeatedly drawing balls without replacement and placing them into $m$ vectors $y^{(1)},\ldots,y^{(m)}\in\{0,1,\star\}^{n}$ that are initially empty until all the vectors are determined. Those vectors in which ($(\neg v_{i})$-ball is placed for some $i\in\{1,\ldots,k\}$ can be ignored, since they will not be accepted by the path $P$ and counted in the weight $\operatorname{wt}(f|_{P})$. Therefore, the above process samples $\operatorname{wt}(f|_{P})$ under the uniform distribution over $f\in\mathcal{B}_{n,m}$.
Denote by $t_{i}$ the random variable that indicates the number of the remaining vectors in the $i$-th step for $i=1,2,\ldots,k$. If the path $P$ computes $f$ (i.e., $f|_{P}\equiv c$), there should be either no vector $(t_{k}=0)$ or full of vectors ($t_{k}=2^{n-k}$) at the end; otherwise, it outputs a wrong answer. So it takes $\Omega(\log m)$ steps to discard all the $m$ vectors if $t_{i}\approx\frac{1}{2}t_{i-1}$ for all $i=1,2,\ldots,\Omega(\log m)$. Since, in each step, $t_{i}\approx\frac{1}{2}t_{i-1}$ holds with high probability by Hoeffding’s inequality, it is unlikely that a short path computes $f$. The above sketch is the proof strategy for Theorem 1.2.
Definition 4.6 ($\delta$-parity path).
[…]
$\frac{1}{2}(1-\delta)\leq\frac{\operatorname{wt}(f|_{\rho_{j}})}{\operatorname{wt}(f|_{\rho_{j-1}})}\leq\frac{1}{2}(1+\delta)$
for every $j=1,2,\ldots,k$.
Lemma 4.7.
Let $P$ be a decision tree path of length at most $\epsilon\log m$. For any $\delta\in(0,\frac{1}{2\epsilon\log m}]$ and $\epsilon\in(0,1)$,
$\Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity for }f\right]<(2\epsilon\log m)\cdot\exp\left(-\frac{1}{2}\cdot\delta^{2}m^{1-\epsilon}\right).$
Proof.
Let $b\leq\epsilon\log m$ denote the length of $P$. The random variable $t_{j}=\operatorname{wt}(f|_{\rho_{j}})$, where $f$ is sampled from $\mathcal{B}_{n,m}$ uniformly at random. Let $X_{j,1},\ldots,X_{j,t_{j-1}}\in\{0,1\}$ be the random variables indicating whether each of the $t_{j-1}$ vectors remains in the $j$-th step, that is, whether it is in the on-set of $f|_{\rho_{j}}$. Random variables $X_{j,1},\ldots,X_{j,t_{j-1}}\in\{0,1\}$ are obtained from sampling without replacement. We have $t_{j}=X_{j,1}+\cdots+X_{j,t_{j-1}}$ and $\mathbb{E}[t_{j}]=\frac{1}{2}t_{j-1}$.
Let $\alpha=\frac{1}{2}(1-\delta)$ and $\beta=\frac{1}{2}(1+\delta)$. The probability that $P$ is not $\delta$-parity for $f$ is exactly
$\ \Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity for }f\right]$
$=$ $\ 1-\Pr\left[\bigwedge_{j=1}^{b}t_{j}\in\left[\alpha t_{j-1},\beta t_{j-1}\right]\right]$
$=$ $\ \Pr\left[\exists 1\leq j\leq b\,\text{ s.t. }\,t_{j}\not\in\left[\alpha t_{j-1},\beta t_{j-1}\right]\wedge\left(\bigwedge_{k=1}^{j-1}t_{k}\in\left[\alpha t_{k-1},\beta t_{k-1}\right]\right)\right].$ (6)
Let $A_{j}$ denote the event $t_{j}\in\left[\alpha t_{j-1},\beta t_{j-1}\right]$ and $B_{j}$ denote the event $t_{j}\in\left[\alpha^{j}m,\beta^{j}m\right]$. By the union bound, (6) is at most
$\sum_{j=1}^{b}\Pr\left[\neg A_{j}\wedge\left(\bigwedge_{k=1}^{j-1}A_{k}\right)\right]\leq\sum_{j=1}^{b}\Pr\left[\neg A_{j}\wedge B_{j-1}\right].$ (7)
If the event $A_{j}$ does not occur, we have $t_{j}>\beta t_{j-1}=\frac{1}{2}(1+\delta)t_{j-1}$ or $t_{j}<\alpha t_{j-1}=\frac{1}{2}(1-\delta)t_{j-1}$, which implies $|t_{j}-\mathbb{E}[t_{j}]|>\frac{1}{2}\delta t_{j-1}$. By Hoeffding’s inequality (Lemma 2.1),
$\Pr\left[\neg A_{j}\wedge B_{j-1}\right]\leq 2\exp\left(-\frac{1}{2}\delta^{2}t_{j-1}\right)\leq 2\exp\left(-\frac{1}{2}\alpha^{j-1}\delta^{2}m\right),$ (8)
since $t_{j-1}\geq\alpha^{j-1}m$ by $B_{j-1}$.
Finally, plugging (8) into (7), we have
$\ \Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity for }f\right]$
$\leq$ $\ \sum_{j=1}^{b}2\exp\left(-\frac{1}{2}\cdot\alpha^{j-1}\delta^{2}m\right)$
$\leq$ $\ 2b\cdot\exp\left(-\frac{1}{2}\cdot\alpha^{b-1}\delta^{2}m\right)$ $\left(\alpha=\frac{1}{2}(1-\delta)<1\right)$
$\leq$ $\ 2b\cdot\exp\left(-\frac{1}{2}\cdot\frac{1}{2^{b-1}}\left(1-\frac{1}{2\epsilon\log m}\right)^{\epsilon\log m-1}\delta^{2}m\right)$ $\left(\delta\leq\frac{1}{2\epsilon\log m}\right)$
$\leq$ $\ 2b\cdot\exp\left(-\frac{1}{2}\delta^{2}\cdot\frac{m}{2^{b}}\right)$ $\left(\left(1-\frac{1}{2x}\right)^{x-1}>\frac{1}{2}\right)$
$\leq$ $\ 2\epsilon\log m\cdot\exp\left(-\frac{1}{2}\cdot\delta^{2}m^{1-\epsilon}\right).$ $\Big{(}b\leq\epsilon\log m\Big{)}$
Definition 4.8 ($(t,\delta)$-parity function).
A boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ is called $(t,\delta)$-parity if any decision tree path of length at most $t$ is a $\delta$-parity path (with respect to $f$).
Lemma 4.9.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a $(t,\delta)$-parity function, and $\operatorname{wt}(f)\leq 2^{n-1}$. If $11\leq t\leq\log\operatorname{wt}(f)-$ 1 and $0\leq\delta\leq\frac{1}{2t}$, then $\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)\geq t$.
Proof.
[…]
$\frac{1}{2^{j}}(1-\delta)^{j}\leq\frac{\operatorname{wt}(f|_{\rho_{j})}}{\operatorname{wt}(f)}=\frac{\operatorname{wt}(f|_{\rho_{1}})}{\operatorname{wt}(f)}\cdot\frac{\operatorname{wt}(f|_{\rho_{2}})}{\operatorname{wt}(f|_{\rho_{1}})}\cdots\frac{\operatorname{wt}(f|_{\rho_{j}})}{\operatorname{wt}(f|_{\rho_{j-1}})}\leq\frac{1}{2^{j}}(1+\delta)^{j}$ (9)
for all $j=1,2,\ldots,k$, where $\rho_{j}$ is the restriction that fixes $x_{i_{p}}$ to $v_{p}$ for $p=1,2,\ldots,j$. Thus, we have $\frac{1}{2^{j}}(1-\delta)^{j}\operatorname{wt}(f)\leq\operatorname{wt}(f|_{\rho_{j}})\leq\frac{1}{2^{j}}(1+\delta)^{j}\operatorname{wt}(f)$. Note that (9) holds for any decision tree path of length at most $t$. So we have
$\frac{1}{2^{j}}(1-\delta)^{j}\operatorname{wt}(f)\leq\operatorname{wt}(f|_{\rho})\leq\frac{1}{2^{j}}(1+\delta)^{j}\operatorname{wt}(f)$ (10)
for any restriction $\rho$ fixing at most $t$ variables.
On one hand, by (10), we have
$\operatorname{wt}(f|_{\rho})$ $\leq(1+\delta)^{t}2^{n-j-1}$ $\Big{(}\operatorname{wt}(f)\leq 2^{n-1}\Big{)}$
$\leq\frac{1}{2}\left(1+\frac{1}{2t}\right)^{t}2^{n-j}$ $\left(\delta\leq\frac{1}{2t}\right)$
$\leq\frac{\sqrt{e}}{2}\cdot 2^{n-j}\leq 0.8244\cdot 2^{n-j}.$ $\left(\left(1+\frac{1}{2n}\right)^{n}\leq\sqrt{e}\right)$
On the other hand, by (10), we have
$\operatorname{wt}(f|_{\rho})$ $\geq(1-\delta)^{t}2^{t-j+1}$ $\Big{(}t\leq\log\operatorname{wt}(f)-1\Big{)}$
$\geq 2\left(1-\frac{1}{2t}\right)^{t}$ $\left(\delta\leq\frac{1}{2t}\text{ and }j\leq t\right)$
$\geq 1.$ $\left(\left(1-\frac{1}{2x}\right)^{x}\geq\frac{1}{2}\text{ when }x\geq 1\right)$
Thus, $\operatorname{wt}(f|_{\rho})$ is less than $2^{n-j}$ and larger than $1$, which implies $f|_{\rho}$ cannot be constant for any restriction $\rho$ fixing at most $t$ variables. Therefore, we conclude the smallest certificate has size $t$ at least, i.e., $\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)\geq t$. ∎
Now, we are ready to prove Theorem 1.2.
Proof of Theorem 1.2.
Let $m=m(n)$ and
$\epsilon=1-\frac{1}{\log m}\left(\log\log n+3\log\log\frac{m}{\log n}+6\right).$
It is straightforward to prove by definition that $\mathrm{cost}(T,x)\geq\mathrm{C}_{x}(f)$ for any input $x\in\{0,1\}^{n}$ and any decision tree $T$ that computes $f$, which implies $\mathrm{D}_{\mathrm{ave}}(f)\geq\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)$. Our goal is to prove
$\Pr_{{f}\sim\mathcal{B}_{n,m}}\left[\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)<\epsilon\log m=\log\frac{m}{\log n}-3\log\log\frac{m}{\log n}-5\right]\to 0$
as $m\rightarrow\infty$. Since $4\log n\leq m\leq 2^{n-1}$, $m,n$ tend to infinity simultaneously,
Let $t=\epsilon\log m$ and $\delta=\frac{1}{2\log\frac{m}{\log n}}\leq\frac{1}{2\epsilon\log m}=\frac{1}{2t}$. Let $\mathrm{len}(P)$ denote the length of a tree path $P$. By Lemma 4.9, if $\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)<t$, then $f$ is not $(b,\delta)$-parity. That is, there exists a path $P$ which is not $\delta$-parity. Thus,
$\ \Pr_{{f}\sim\mathcal{B}_{n,m}}\left[\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)<\epsilon\log m\right]$
[…]
By the union bound, (11) is at most
$\ \sum_{\mathrm{len}(P)<t}\Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity for }f\right]$
[…]
$\leq$ $\ n^{\epsilon\log m}(\epsilon\log m)^{2\epsilon\log m}\cdot(2\epsilon\log m)\cdot\exp\left(-\frac{1}{2}\delta^{2}m^{1-\epsilon}\right)$
$\leq$ $\ \exp\left(\ln n\cdot 4\epsilon\log m\right)\cdot\exp\left(-\frac{m^{1-\epsilon}}{8(\log\frac{m}{\log n})^{2}}\right).$ (12)
Then, since $\epsilon\log m\leq\log\frac{m}{\log n}$, (12) is at most
$\ \exp\left(\log n\cdot(4\ln 2)\cdot\log\frac{m}{\log n}-\frac{m^{1-\epsilon}}{8(\log\frac{m}{\log n})^{2}}\right)$
$=$ $\ \exp\left(-4(1-\ln 2)\log\frac{m}{\log n}\cdot\log n\right)$
$\leq$ $\ \exp\left(-8(1-\ln 2)\log n\right),$ $\left(m\geq 4\log n\right)$
which tends to zero as $n\rightarrow\infty$. Therefore, we conclude
$\mathrm{D}_{\mathrm{ave}}(f)\geq\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)\geq\epsilon\log m=\log\frac{m}{\log n}-3\log\log\frac{m}{\log n}-5$
for almost all functions $f\in\mathcal{B}_{n,m}$. That is, $\mathrm{D}_{\mathrm{ave}}(f)$ is at least $\log\frac{m}{\log n}-O(\log\log\frac{m}{\log n})$, since $m\geq 4\log n$. ∎
5 Circuits
In this section, we study $\mathrm{D}_{\mathrm{ave}}(f)$ of circuits that consist of AND, OR, NOT gates with unbounded fan-in. Recall that the size of a circuit is defined as the number of the gates and that the depth is the length of the longest path between sink and source.
As a warm-up, we show an upper bound of $\mathrm{D}_{\mathrm{ave}}(F)=O(s)$ for general size-$s$ circuits $F$. The bound is tight up to a multiplicative factor, since $\mathrm{D}_{\mathrm{ave}}(\mathrm{XOR}_{n})=n$, and the function $\mathrm{XOR}_{n}$ is computable by a circuit of size $O(n)$ and depth $O(\log n)$.
Lemma 5.1.
$\mathrm{D}_{\mathrm{ave}}(F)\leq 2s$ for every circuit $F$ of size $s$.
Proof.
Since the average cost of evaluating each AND/OR/NOT gate does not exceed $2$ (Corollary B.2), it costs at most $2s$ on average to evaluate all $s$ gates. ∎
Let us define random restriction and criticality.
Definition 5.2.
A $p$-random restriction, denoted by $\mathcal{R}_{p}$, is a distribution over restrictions leaving $x_{i}$ unset with probability $p$ and fixing $x_{i}$ to 00 or $1$ with equal probability $\frac{1}{2}(1-p)$ independently for each $i=1,2,\ldots,n$.
Definition 5.3 ([Ros17, Ros19]).
A boolean function $f$ is $\lambda$-critical if and only if
$\Pr_{\rho\sim\mathcal{R}_{p}}\left[\mathrm{D}(f|_{\rho})\geq t\right]\leq(p\lambda)^{t}$
for every $p\in[0,1]$ and $t\in\mathbb{N}$.
The next lemma gives an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$ for any $\lambda$-critical function $f$.
Lemma 5.4.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be $\lambda$-critical. Then, for any $\epsilon>0$,
$\mathrm{D}_{\mathrm{ave}}(f)\leq n\left(1-\frac{1}{(1+\epsilon)\lambda}\right)+\frac{1+\epsilon}{\epsilon^{2}}.$ (13)
Proof.
Let $p=\frac{1}{(1+\epsilon)\lambda}$. Since $f$ is $\lambda$-critical, we have $\Pr_{\rho\sim\mathcal{R}_{p}}[\mathrm{D}(f|_{\rho})\geq t]\leq(1+\epsilon)^{-t}$. Consider a query algorithm that queries each variable independently with probability $1-p$, and then applies a worst-case optimal query algorithm to the remaining subfunction $f|_{\rho}$. We have
$\mathrm{D}_{\mathrm{ave}}(f)$ $\leq\mathop{\mathbb{E}}_{\rho\sim\mathcal{R}_{p}}[|\operatorname{supp}(\rho)|+\mathrm{D}_{\mathrm{ave}}(f|_{\rho})]$
$\leq n(1-p)+\sum_{t=0}^{\infty}t\cdot\Pr_{\rho\sim\mathcal{R}_{p}}[\mathrm{D}(f|_{\rho})=t]$
$\leq n\left(1-p\right)+\sum_{t=0}^{\infty}\frac{t}{(1+\epsilon)^{t}}$
$=n\left(1-\frac{1}{(1+\epsilon)\lambda}\right)+\frac{1+\epsilon}{\epsilon^{2}}.$
Remark 5.5.
When $\epsilon=1$, one can give an alternative proof of the upper bound by applying the OS inequality $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\mathrm{DT}_{\mathrm{size}}(f)$ (Theorem A.2) and the bound $\mathrm{DT}_{\mathrm{size}}(f)\leq O(2^{n(1-\frac{1}{2\lambda})})$ [Ros19], where $\mathrm{DT}_{\mathrm{size}}(f)$ denotes the decision tree size.
Remark 5.6.
Letting $\epsilon=(\frac{\lambda}{n})^{\frac{1}{3}}$, we have $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{\lambda})+\Theta\left(\frac{n}{\lambda}\right)^{\frac{2}{3}}$ for $\lambda$-critical functions $f$.
Combining Lemma 5.4 with the previous results of criticality, one can derive upper bounds on $\mathrm{D}_{\mathrm{ave}}(f)$ for various circuit/formula classes. These bounds are incomparable with Lemma 5.1.
Corollary 5.7 ([Has86]).
For every CNF or DNF formula $F:\{0,1\}^{n}\rightarrow\{0,1\}$ of width $w$,
$\mathrm{D}_{\mathrm{ave}}(F)\leq n\left(1-\frac{1}{O(w)}\right).$
Corollary 5.8 ([Ros17, Ros19]).
For every CNF or DNF formula $F:\{0,1\}^{n}\rightarrow\{0,1\}$ of size $s$,
$\mathrm{D}_{\mathrm{ave}}(F)\leq n\left(1-\frac{1}{O(\log s)}\right).$
Corollary 5.9 ([Has14, Ros17]).
For every $\mathsf{AC}^{0}$ circuit $F:\{0,1\}^{n}\rightarrow\{0,1\}$ of size $s$ and depth $d$,
$\mathrm{D}_{\mathrm{ave}}(F)\leq n\left(1-\frac{1}{O(\log s)^{d-1}}\right).$
Corollary 5.10 ([HMS23]).
For every $\mathsf{AC}^{0}$ formula $F:\{0,1\}^{n}\rightarrow\{0,1\}$ of size $s$ and depth $d$,
$\mathrm{D}_{\mathrm{ave}}(F)\leq n\left(1-\frac{1}{O(\frac{1}{d}\log s)^{d-1}}\right).$
As we have discussed in the introduction, it is interesting to prove or disprove the tightness of these upper bounds. In order to provide evidence on the tightness of Corollary 5.7, 5.8, we prove Theorem 1.3 and show a DNF formula of width $w$ and size $2^{w}/w$ with $\mathrm{D}_{\mathrm{ave}}(f)=n\left(1-\frac{\log n}{\Theta(w)}\right)$.
Here we briefly explain how to construct the DNF formula. Although $\mathrm{D}_{\mathrm{ave}}(\mathrm{OR}_{n})=O(1)$ under a uniform input distribution, it costs $n(1-o(1))$ on average to determine the output of the OR function under a $p$-biased input distribution when $p=o(1/n)$ (Exercise 8.65 in [ODo14]). Our approach is to employ a biased function $g$ with $p=\Pr_{x\in\{0,1\}^{n}}[g(x)=1]$ and $\mathrm{D}_{\mathrm{ave}}(g)=n(1-o(1))$, which is given by Theorem 1.2, as a “simulator” of $p$-biased variable. Then, we show the composition $\mathrm{OR}\circ g$ is hard to query under a uniform distribution and is computable by a somewhat small DNF formula. As such, Theorem 1.3 follows.
Proof of Theorem 1.3.
Let $s=\lceil\frac{2^{w}}{w}\rceil$ and $h=\lceil\frac{n}{w}\rceil$. Define $m=\lceil\frac{s}{h}\rceil$. Since $w\leq n$, we have
$m\leq\frac{s}{h}+1\leq\frac{\frac{2^{w}}{w}+1}{\frac{n}{w}}+1=\frac{2^{w}}{n}+\frac{w}{n}+1\leq\frac{2^{n}}{n}+2\leq 2^{n-1}$
for any $n\geq 4$. Since $w\geq\log n+\log\log n+3$, we have $m\geq 4\log n$. By Theorem 1.2, there exists $g\in\mathcal{B}_{w,m}$ such that
$d=\min_{y\in\{0,1\}^{w}}\mathrm{C}_{y}(g)\geq\log\frac{m}{\log w}-O\left(\log\log\frac{m}{\log w}\right)=w-\log n-O(\log w).$
Let $p=\frac{m}{2^{w}}$ denote the probability that the function $g$ outputs $1$. Note that $p\leq\frac{1}{n}+\frac{w}{n\cdot 2^{w}}+\frac{1}{2^{w}}\leq\frac{3}{n}$.
Let
$f(x)=\bigvee_{k=1}^{h}g(x^{(k)}),$ (14)
where $x=(x^{(1)},\ldots,x^{(h)})\in\{0,1\}^{n}$ and $x^{(1)},\ldots,x^{(h)}\in\{0,1\}^{w}$. It is obvious that $f$ is computable by a DNF formula of width $w$ and size $mh\leq s$, because each individual $g$ is computable by a DNF formula of width $w$ and size $m$.
Let $T$ be any query algorithm computing $f$. Let us condition on the event $g(x^{(1)})=\cdots=g(x^{(h)})=0$, which happens with probability $(1-p)^{h}$. The algorithm $T$ needs to query at least $d$ variables to evaluate each clause $g(x^{(k)})$. If $g(x^{(1)})=\cdots=g(x^{(h)})=0$, then $T$ queries at least $hd$ variables. Thus,
$\mathrm{D}_{\mathrm{ave}}(f)\geq$ $\ hd\cdot\Pr_{x\in\{0,1\}^{n}}\left[g(x^{(1)})=\cdots=g(x^{(h)})=0\right]$
$\geq$ $\ n\left(1-\frac{\log n+O(\log w)}{w}\right)(1-p)^{h}$ $\Big{(}d\geq w-\log n-O(\log w)\Big{)}$
$\geq$ $\ n\left(1-\frac{O(\log n)}{w}\right)(1-ph)$ $\Big{(}1-ph\leq(1-p)^{h}\Big{)}$
$=$ $\ n\left(1-\frac{\log n}{\Omega(w)}\right).$ $\left(ph\leq\frac{3}{w}\left(1+\frac{w}{n}\right)\right)$ (15)
On the other hand, we can query $f$ by evaluating all the clauses one by one. By Lemma 4.1, $\mathrm{D}_{\mathrm{ave}}(g)\leq\log m+2\leq w-\log n+4$. Thus,
$\mathrm{D}_{\mathrm{ave}}(f)\leq h\cdot\mathrm{D}_{\mathrm{ave}}(g)\leq n\left(1-\frac{\log n-4}{w}\right)=n\left(1-\frac{\log n}{O(w)}\right).$ (16)
Finally, combining (15) and (16), we conclude $\mathrm{D}_{\mathrm{ave}}(f)=n\left(1-\frac{\log n}{\Theta(w)}\right)$. ∎
6 Conclusion
In this paper, we study the average-case query complexity of boolean functions under the uniform distribution. We give an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$ in terms of its weight. We prove, for almost all fixed-weight boolean functions, our upper bound is tight up to an additive logarithmic term). We show that, for any $w\geq\log n+\sqrt{\log n}$, there exists a DNF formula of width $w$ and size $2^{w}/w$ such that $\mathrm{D}_{\mathrm{ave}}(f)=n(1-\frac{\log n}{\Theta(w)})$, which suggests that the switching lemmas are almost tight. We also study or compute $\mathrm{D}_{\mathrm{ave}}(f)$ of several functions, including the penalty shoot-out function, symmetric functions, linear threshold functions and the tribes functions.
Theorems 1.1 and 1.2 essentially relate $\mathrm{D}_{\mathrm{ave}}(f)$ to the zero-order Fourier coefficient $\widehat{f}(\{\emptyset\})$. Establishing an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$ in terms of higher-order Fourier coefficients (e.g. influences) can deepen our understanding of $\mathrm{D}_{\mathrm{ave}}(f)$ and pave the way for more applications. For example, it is unknown whether the lower bound $\mathrm{D}_{\mathrm{ave}}(\mathrm{Hex}_{L\times L})\geq L^{1.5+o(1)}$, which coincides with the expected optimal playing time for the random-turn Hex game, is tight or not [PSS+07], and such a bound on $\mathrm{D}_{\mathrm{ave}}(f)$ might shed light on the question.
It remains open to prove the tightness of the upper bounds (Corollary 5.7, 5.8) for CNF/DNF formulas. If true, it would imply that the switching lemmas are tight.
Acknowledgements
We are grateful to the anonymous reviewers for their valuable feedback.
[…]
Appendix A General lower bounds
Theorem A.1 (OSSS inequality [OSS+05]).
For any decision tree $T$ computing a boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$,
$\sum_{i=1}^{n}\delta_{i}\cdot\mathrm{{Inf}}_{i}[f]\geq\mathrm{{Var}}[f],$
where $\delta_{i}\in[0,1]$ is the probability that the $i$-th variable $x_{i}$ is queried by the decision tree $T$. In particular, taking $T$ as an optimal tree computing $f$ in terms of $\mathrm{D}_{\mathrm{ave}}(f)$, we have
$\mathrm{D}_{\mathrm{ave}}(f)=\sum_{i=1}^{n}\delta_{i}\geq\frac{\mathrm{{Var}}[f]}{\max_{i\in[n]}\mathrm{{Inf}}_{i}[f]}.$
Theorem A.2 (OS inequality [OS06]).
For every boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$,
$\left(\sum_{i=1}^{n}\widehat{f}(\{i\})\right)^{2}\leq\mathrm{D}_{\mathrm{ave}}(f)\leq\log\mathrm{DT}_{\mathrm{size}}(f),$
where $\mathrm{DT}_{\mathrm{size}}(f)$ is the decision tree size. In particular, $\mathrm{D}_{\mathrm{ave}}(f)\geq\mathrm{s}_{\mathrm{ave}}(f)^{2}$ when $f$ is monotone.
Appendix B Special function classes
B.1 AND, OR and XOR
Define $\mathrm{AND}_{n}(x)=\bigwedge_{i=1}^{n}x_{i}$ and $\mathrm{OR}_{n}(x)=\bigvee_{i=1}^{n}x_{i}$ and $\mathrm{XOR}_{n}(x)=\bigoplus_{i=1}^{n}x_{i}$. The following proposition gives the exact value of $\mathrm{D}_{\mathrm{ave}}(f)$ for boolean functions $f$ with weight $1$, including the AND function. It shows a super-exponential gap between $\mathrm{D}(f)$ and $\mathrm{D}_{\mathrm{ave}}(f)$, since $\mathrm{D}(\mathrm{AND}_{n})=n$ and $\mathrm{D}_{\mathrm{ave}}(\mathrm{AND}_{n})=O(1)$, and also gives the exact value of $\mathrm{D}_{\mathrm{ave}}(\mathrm{OR}_{n})$.
Proposition B.1.
$\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})$ for any $n$-variable boolean function $f$ with $\operatorname{wt}(f)=1$.
Proof.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function with a unique black point $z\in\{0,1\}^{n}$. We prove $\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})$ by induction on $n$. When $n=1$, it is easy to verify $\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})=1$.
Suppose an optimal query algorithm queries $x_{i}$ first. If $x_{i}\not=z_{i}$, the algorithm outputs 00. Otherwise, the algorithm continues on the subfunction $f|_{x_{i}=z_{i}}$. Therefore,
$\mathrm{D}_{\mathrm{ave}}(f)$ $=1+\Pr_{x\in\{0,1\}^{n}}\left[x_{i}\neq z_{i}\right]\cdot 0+\Pr_{x\in\{0,1\}^{n}}\left[x_{i}=z_{i}\right]\cdot\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=z_{i}})$
$=1+\frac{1}{2}\cdot 2\left(1-\frac{1}{2^{n-1}}\right)$
$=2\left(1-\frac{1}{2^{n}}\right),$
where the second step is by the induction hypothesis. ∎
Corollary B.2 (Exercise 8.65 in [ODo14]).
$\mathrm{D}_{\mathrm{ave}}(\mathrm{AND}_{n})=\mathrm{D}_{\mathrm{ave}}(\mathrm{OR}_{n})=2(1-\frac{1}{2^{n}})$.
Proof.
By Proposition 3.6, $\mathrm{D}_{\mathrm{ave}}(\mathrm{AND}_{n})=\mathrm{D}_{\mathrm{ave}}(\mathrm{OR}_{n})$, since the OR function is the dual of the AND function. Then, the corollary follows from Proposition B.1. ∎
In addition, we show the XOR function (and its negation) is the only function with the maximum complexity $\mathrm{D}_{\mathrm{ave}}(f)=n$.
Proposition B.3.
Let $f$ be an $n$-variable boolean function. Then, $\mathrm{D}_{\mathrm{ave}}(f)=n$ if and only if $f=\mathrm{XOR}_{n}$ or $f=\mathrm{XOR}_{n}\oplus 1$.
Proof.
If $f$ is $\mathrm{XOR}_{n}$ or $\mathrm{XOR}_{n}\oplus 1$, then $\mathrm{D}_{\mathrm{ave}}(f)=n$, since $f|_{\rho}$ is not constant for any restriction $\rho$ fixing $\leq n-1$ variables.
Now we use induction to prove the opposite direction that if $\mathrm{D}_{\mathrm{ave}}(f)=n$, then $f$ must be $\mathrm{XOR}_{n}$ or $\mathrm{XOR}_{n}\oplus 1$. When $n=1$, the conclusion is obvious. Assuming the conclusion is true for $n-1$, we prove it for $n$. Suppose an optimal query algorithm computing $f$ is to query the variable $x_{i}$ first. Since $\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=b})\leq n-1$ for any $b\in\{0,1\}$, we have
$n=\mathrm{D}_{\mathrm{ave}}(f)\leq 1+\frac{1}{2}\left(\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=0})+\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=1})\right)\leq n,$
which implies that $\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=0})=\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=1})=n-1$. By the induction hypothesis, for any $b\in\{0,1\}$, we know $f|_{x_{i}=b}$ is either $\mathrm{XOR}_{n-1}\text{ or }\mathrm{XOR}_{n-1}\oplus 1$.
If $f|_{x_{i}=0}=f|_{x_{i}=1}$, meaning $f=f|_{x_{i}=0}$, then the variable $x_{i}$ is not pivotal on any input, which contradicts our assumption that $\mathrm{D}_{\mathrm{ave}}(f)=n$. Thus, $f|_{x_{i}=0}$ and $f|_{x_{i}=1}$ are not the same. We conclude that $f=\mathrm{XOR}_{n}$ if $f|_{x_{i}=0}=\mathrm{XOR}_{n-1}$, and that $f=\mathrm{XOR}_{n}\oplus 1$ if $f|_{x_{i}=0}=\mathrm{XOR}_{n-1}\oplus 1$. ∎
B.2 Penalty shoot-out function
Besides the AND/OR functions, which are highly biased, there are monotone balanced functions such that the gap between $\mathrm{D}_{\mathrm{ave}}(f)$ and $\mathrm{D}(f)$ is arbitrarily large.
Let us define the penalty shoot-out function $\mathrm{PSO}_{n}:\{0,1\}^{2n+1}\rightarrow\{0,1\}$ as follows. Consider an $n$-round penalty shoot-out in a football game. In each round, two teams, A and B, each take a penalty kick in turn, with team A going first. Let $x_{2i-1}=1$ indicate the event team A scores in the $i$-th round, and let $x_{2i}=0$ — this is to make the function monotone — indicate the event that team B scores in the $i$-th round for $i=1,2,\ldots,n$. The game continues until one team scores and the other does not (within the same round). If no winner is declared after $2n$ kicks, an additional kick by team A decides the game. In this final kick, if team A scores, team A wins and $\mathrm{PSO}_{n}(x)=1$; otherwise, team B wins and $\mathrm{PSO}_{n}(x)=0$. To the best of our knowledge, $\mathrm{PSO}_{n}$ is first studied here.
Assume both teams have equal probabilities of scoring, that is, $\mathrm{PSO}_{n}$ is defined under a uniform distribution. The function $\mathrm{PSO}_{n}$ is a monotone balanced function with $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=O(1)$ and $\mathrm{D}(\mathrm{PSO}_{n})=\Theta(n)$.
Proposition B.4.
$\mathrm{D}(\mathrm{PSO}_{n})=2n+1$ and $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=4-\frac{3}{2^{n}}$.
Proof.
In the worst case, the winner cannot be declared until the last round is finished. So $\mathrm{D}(\mathrm{PSO}_{n})=2n+1$.
In the average case, we prove $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=4-\frac{3}{2^{n}}$ by induction on $n$. When $n=0$, $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=1$, because one kicked by team A decides the game. Assuming $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n-1})=4-\frac{3}{2^{n-1}}$ holds, let us prove $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=4-\frac{3}{2^{n}}$. In each round, there are four cases:
  • (1)
    Team A scores, and team B does not.
  • (2)
    Team B scores, and team A does not.
  • (3)
    Both teams score.
  • (4)
    Neither scores.
Each case happens with equal probability $\frac{1}{4}$. If case 1) or 2) happens, the winner is decided; if case 3) or 4) happens, the game will continue. Therefore, we have
$\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})$ $=\frac{1}{2}\cdot 2+\frac{1}{2}\cdot\left(\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n-1})+2\right)$
$=\frac{1}{2}\left(4-\frac{3}{2^{n-1}}\right)+2$
$=4-\frac{3}{2^{n}},$
completing the induction proof. ∎
B.3 Symmetric functions
Definition B.5.
Say a function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ is symmetric if and only if $f(x)=f(x^{\pi})$ for every permutation $\pi\in S_{n}$, where $(x^{\pi})_{i}=x_{\pi(i)}$ for all $i=1,2,\ldots,n$.
Symmetric functions are useful and contain many crucial examples, such as the threshold function, which was used to show the super-exponential gap between $\mathrm{D}_{\mathrm{ave}}(f)$ and $\mathrm{R}_{\mathrm{ave}}(f)$.
Definition B.6.
The threshold function is defined by $\mathrm{Thr}_{n,t}(x)=1$ if and only if $|x|\geq t$.
Proposition B.7 (Theorem 3.1 in [AD01]).
$\mathrm{R}_{\mathrm{ave}}(\mathrm{Thr}_{n,\frac{n}{10}})=O(1)$ and $\mathrm{D}_{\mathrm{ave}}(\mathrm{Thr}_{n,\frac{n}{10}})=\Omega(n)$.
The value $f(x)$ only depends on input’s weight $|x|$ for any symmetric function $f$, which can be viewed as a function mapping $\{0,\ldots,n\}$ to $\{0,1\}$. For convenience, we say a symmetric function $f$ is constant on an interval $S=\{a,\ldots,b\}\subseteq\{0,\ldots,n\}$ if $f(x)$ is constant for all inputs $x$ with weight $|x|\in S$. An optimal query algorithm computing $f$ is to query variables arbitrarily until the remaining subfunction is constant on some interval. Based on this observation, the following lemmas give exact characterizations of $\mathrm{D}_{\mathrm{ave}}(f)$ for symmetric functions.
Lemma B.8.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be symmetric. Let $I_{0}=\{b_{0}+1,\ldots,b_{1}\},\ldots,I_{m}=\{b_{m}+1,\ldots,b_{m+1}\},$ where $-1=b_{0}<b_{1}<\cdots<b_{m}<b_{m+1}=m$, be all the maximal intervals such that $f$ becomes constant when input’s weight is restricted to each interval. Then, we have
$\mathrm{D}_{\mathrm{ave}}(f)=\frac{1}{2^{n}}\sum_{i=0}^{m}\sum_{k=0}^{a_{i}-1}\left({n-k-1\choose b_{i}}+{n-k-1\choose n-b_{i+1}-1}\right)2^{k}(n-k),$
where $a_{i}=b_{i+1}-b_{i}$ for $i=0,1,\ldots,m$.
Proof.
Consider $I_{i}=\{b_{i}+1,\ldots,b_{i+1}\}$ for an example. Assume we query $n-k$ variables and get $p$ ones and $q$ zeros. Observe that $p\leq|x|\leq n-q$, since $x$ has contained $p$ ones and $q$ zeros. In the best case (i.e., $p=b_{i}+1$ and $n-q=b_{i+1}$), we can determine that $|x|\in I_{i}$ by querying only $(b_{i}+1)+(n-b_{i+1})=n-a_{i}+1$ variables. In the worst case, we would need to query all $n$ variables. Thus, $k$ ranges from 00 to $a_{i}-1$.
Since $p\leq|x|\leq n-q$, we have $|x|\in I_{i}$ if and only if $p\geq b_{i}+1$ and $q\geq n-b_{i+1}$. Since both conditions cannot be satisfied at the same time, we can convince $|x|\in I_{i}$ and stop querying as long as (1) $p=b_{i}+1$ and $q\geq n-b_{i+1}$; or (2) $p\geq b_{i}+1$ and $q=n-b_{i+1}$.
In the case that $p=b_{i}+1$ and $q\geq n-b_{i+1}$, the last queried variable must be $1$, and the previous $n-k-1$ bits must contain $b_{i}$ ones. Therefore, the probability of querying only $n-k$ variables is exactly $\frac{1}{2^{n-k}}{n-k-1\choose b_{i}}.$ A similar argument works for the case that $p\geq b_{i}+1$ and $q=n-b_{i+1}$. As such, we derive the desired formula. ∎
Proposition B.9.
Let $\mathcal{R}_{k}$ denote the uniform distribution over restrictions which fix exactly $k$ variables. For any symmetric function $f:\{0,1\}^{n}\rightarrow\{0,1\}$, we have
$\mathrm{D}_{\mathrm{ave}}(f)=\sum_{k=0}^{n-1}\Pr_{\rho\sim\mathcal{R}_{k}}\left[f|_{\rho}\not\equiv\mathrm{const}\right].$
Proof.
Since $f$ is symmetric, the order in which the variables are queried does not matter. Thus, an optimal strategy is to query $x_{1},x_{2},\ldots,x_{n}$ in order, until the function is a constant; Denote the corresponding decision tree by $T$. We have
$\mathrm{D}_{\mathrm{ave}}(f)$ $=\mathbb{E}_{x\in\{0,1\}^{n}}[\mathrm{cost}(T,x)]$
$=\sum_{k=1}^{n}\Pr_{x\in\{0,1\}^{n}}\left[\mathrm{cost}(T,x)=k\right]\cdot k$
$=\sum_{k=0}^{n-1}\Pr_{x\in\{0,1\}^{n}}\left[\mathrm{cost}(T,x)>k\right].$ (17)
Note that $\Pr_{x\in\{0,1\}^{n}}\left[\mathrm{cost}(T,x)>k\right]=\Pr_{\rho\sim\mathcal{R}_{k}}\left[f|_{\rho}\not\equiv\mathrm{const}\right]$, which completes the proof. ∎
Definition B.10.
The majority function is defined by $\mathrm{Maj}_{n}=\mathrm{Thr}_{n,\lceil\frac{n+1}{2}\rceil}$.
The majority function represents a symmetric function that is constant on the sets $S_{0}=\{0,\ldots,\lfloor n/2\rfloor\}$ and $S_{1}=\{\lfloor n/2\rfloor+1,\ldots,n\}$. Combining Lemma B.8 with basic identities of binomial coefficients, we can calculate the exact value of $\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{n})$.
Lemma B.11 ([Len93]).
$\sum_{d=0}^{k}{k+d\choose d}\frac{1}{2^{d}}=2^{k}$ for every $k\in\mathbb{N}$.
Proposition B.12.
$\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{n})=n-\Theta(\sqrt{n})$.
Proof.
Observe that $\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{n})=1+\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{n-1})$ when the integer $n$ is odd. All we have to do is to prove $\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{n})=n-\Theta(\sqrt{n})$ for $n\in 2\mathbb{N}$.
Let $n=2k$. The majority function is the symmetric function which is constant on the sets $\{0,\ldots,k\}$ and $\{k+1,\ldots,2k\}$. By Lemma B.8,
$\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{n})$ $=\frac{1}{2^{n}}\sum_{d=0}^{k}\left[{2k-d-1\choose k-1}+{2k-d-1\choose k}\right]2^{d}(n-d)$
$=\frac{1}{2^{n}}\sum_{d=0}^{k}{2k-d\choose k}2^{d}(n-d)$ (Pascal’s triangle)
$=\frac{1}{2^{k}}\sum_{d=0}^{k}{k+d\choose d}\frac{k+d}{2^{d}}.$
Then, by Lemma B.11, we have
$\frac{1}{2^{k}}\sum_{d=0}^{k}{k+d\choose d}\frac{k+d}{2^{d}}$ $=k+\frac{1}{2^{k}}\sum_{d=1}^{k}{k+d-1\choose d-1}\frac{k+d}{2^{d}}$
$=k+\frac{1}{2}-{2k\choose k}\frac{2k+1}{2^{2k+1}}+\frac{1}{2}\left(\frac{1}{2^{k}}\sum_{d=0}^{k}{k+d\choose d}\frac{k+d}{2^{d}}\right)$
$=2k+1-{2k\choose k}\frac{2k+1}{4^{k}}.$
Finally, since (${2k\choose k}=\frac{4^{k}}{\sqrt{\pi k}}(1-o(1))$, we conclude that $\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{n})=n-\Theta(\sqrt{n})$. ∎
Intuitively, for any symmetric function $f$ that is not constant near the middle of the interval $\{0,\ldots,n\}$, such as the majority function, $\mathrm{D}_{\mathrm{ave}}(f)$ is expected to approach $n(1-o(1))$. The next lemma confirms this hunch.
Lemma B.13.
Let $f:\{0,1\}^{n}\to\{0,1\}$ be a symmetric function. For any constant $c>0$, if $f$ is non-constant on inputs with weight $|x|\in\{\frac{n}{2}-c\sqrt{n},\ldots,\frac{n}{2}+c\sqrt{n}\}$, then $\mathrm{D}_{\mathrm{ave}}(f)\geq n-O_{c}(\sqrt{n})$.
Proof.
For convenience, let $\alpha=\frac{n}{2}-c\sqrt{n}$. By Proposition B.9, we have $\mathrm{D}_{\mathrm{ave}}(f)=\sum_{k=0}^{n-1}\Pr_{\rho\sim\mathcal{R}_{k}}[f|_{\rho}\not\equiv\mathrm{const}]$. Observe that, for every restriction $\rho\in\mathcal{R}_{i}$, the subfunction $f|_{\rho}$ is not constant if $|\rho^{-1}(1)|\leq\alpha$ and $n-|\rho^{-1}(0)|=n-i+|\rho^{-1}(1)|\geq n-\alpha$. So we have
$\sum_{i\geq 1}\Pr_{\rho\in\mathcal{R}_{i}}[f|_{\rho}\not\equiv\mathrm{const}]$ $\geq\sum_{i=1}^{2\alpha}\Pr_{\rho\in\mathcal{R}_{i}}\left[i-\alpha\leq|\rho^{-1}(1)|\leq\alpha\right]$
$=\alpha+\sum_{i=\alpha+1}^{2\alpha}\Pr_{\rho\in\mathcal{R}_{i}}\left[i-\alpha\leq|\rho^{-1}(1)|\leq\alpha\right]$
$=2\alpha-2\sum_{i=\alpha+1}^{2\alpha}\Pr_{\rho\in\mathcal{R}_{i}}\left[|\rho^{-1}(1)|<i-\alpha\right],$ (18)
where the second step is because $|\rho^{-1}(1)|\leq\alpha$ for any $\rho\in\mathcal{R}_{i}$ when $i\leq\alpha$, and the last step is because $\Pr_{\rho\in\mathcal{R}_{i}}\left[i-\alpha\leq|\rho^{-1}(1)|\leq\alpha\right]=1-\Pr_{\rho\in\mathcal{R}_{i}}\left[|\rho^{-1}(1)|<i-\alpha\right]-\Pr_{\rho\in\mathcal{R}_{i}}\left[|\rho^{-1}(1)|>\alpha\right]=1-2\Pr_{\rho\in\mathcal{R}_{i}}\left[|\rho^{-1}(1)|<i-\alpha\right]$.
Let us upper bound the term $2\sum_{i=\alpha+1}^{2\alpha}\Pr_{\rho\in\mathcal{R}_{i}}\left[|\rho^{-1}(1)|<i-\alpha\right]$ in (18). We have
[…]
Plugging (19) into (18), we have
$\mathrm{D}_{\mathrm{ave}}(f)$ $\geq n-2c\sqrt{n}-4\sqrt{\pi\left(\frac{n}{2}-c\sqrt{n}\right)}\geq n-(2c+2\sqrt{2\pi})\sqrt{n}$
as desired. ∎
Generalizing the majority function, we prove $\mathrm{D}_{\mathrm{ave}}(f)\leq n-\Omega_{c}(\sqrt{n})$ for symmetric functions which are constant on the intervals $\{0,\ldots,\frac{n}{2}-c\sqrt{n}-1\}$ and $\{\frac{n}{2}+c\sqrt{n}+1,\ldots,n\}$.
Lemma B.14 ([MV01]).
Let $X_{1},\cdots,X_{n}$ be independent random variables, each attaining the values 00 and $1$, both with probability $\frac{1}{2}$. Let $X=X_{1}+\cdots+X_{n}$. Then, we have, for any integer $t\in\{0,\ldots,\frac{n}{8}\}$,
$\Pr\left[X>\frac{n}{2}+t\right]\geq\frac{1}{27}\exp\left(-\frac{24t^{2}}{n}\right).$
Lemma B.15.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a symmetric function. If there exists a constant $c>0$ such that $f$ is constant on inputs with weight $|x|\in\{0,\ldots,\frac{n}{2}-c\sqrt{n}\}$, or $f$ is constant on inputs with weight $|x|\in\{\frac{n}{2}+c\sqrt{n},\ldots,n\}$, then $\mathrm{D}_{\mathrm{ave}}(f)\leq n-\Omega_{c}(\sqrt{n})$.
Proof.
Let $\beta=\frac{n}{2}+c\sqrt{n}$. By Proposition B.9, we have $\mathrm{D}_{\mathrm{ave}}(f)=\sum_{k=0}^{n-1}\Pr_{\rho\sim\mathcal{R}_{k}}[f|_{\rho}\not\equiv\mathrm{const}]$. When $i>\beta$, the subfunction $f|_{\rho}$ is constant if
  • $|\rho^{-1}(1)|>\beta$, since $f$ is constant on inputs with weight $|x|\in\{\beta,\ldots,n\}$; or
  • $|\rho^{-1}(0)|=i-|\rho^{-1}(1)|>\beta$, since $f$ is constant on inputs with weight $|x|\in\{0,\ldots,n-\beta\}$.
When $i\leq\beta$, the subfunction $f|_{\rho}$ may be non-constant, so we upper bound $\Pr_{\rho\in\mathcal{R}_{i}}[f|_{\rho}\not\equiv\mathrm{const}]$ by 1. We have
$\sum_{i=1}^{n}\Pr_{\rho\in\mathcal{R}_{i}}[f|_{\rho}\not\equiv\mathrm{const}]$ $\leq\beta+\sum_{i=\beta+1}^{n}\left(1-\Pr_{\rho\in\mathcal{R}_{i}}[f|_{\rho}\equiv\mathrm{const}]\right)$
$\leq\beta+\sum_{i=\beta+1}^{n}\left(1-\max\left(\Pr_{\rho\in\mathcal{R}_{i}}\left[|\rho^{-1}(1)|<i-\beta\right],\Pr_{\rho\in\mathcal{R}_{i}}\left[|\rho^{-1}(1)|>\beta\right]\right)\right)$
$=n-\sum_{i=\beta+1}^{n}\Pr_{\rho\in\mathcal{R}_{i}}[|\rho^{-1}(1)|>\beta].$ (20)
The term $\sum_{i=\beta+1}^{n}\Pr_{\rho\in\mathcal{R}_{i}}[|\rho^{-1}(1)|>\beta]$ in (20) is at least
[…]
$\geq\frac{c}{27}e^{-54c^{2}(1+o(1))}\sqrt{n}$ (21)
Plugging (21) into (20), we conclude that $\mathrm{D}_{\mathrm{ave}}(f)\leq n-\Omega_{c}(\sqrt{n})$. ∎
B.4 Linear threshold function
Definition B.16.
A linear threshold function (LTF) $f:\{\pm 1\}^{n}\rightarrow\{\pm 1\}$ is a boolean function that can be represented as $f(x)=\mathrm{sgn}(w_{0}+\sum_{i=1}^{n}w_{i}x_{i})$ for some real numbers $w_{0},\ldots,w_{n}\in\mathbb{R}$ (called weights), where $\mathrm{sgn}(t)=+1$ if $t\geq 0$ and $\mathrm{sgn}(t)=-1$ if $t<0$.
Linear threshold functions can equivalently be defined over the set $\{0,1\}^{n}$, but Definition B.16 is more concise and straightforward. Say a query algorithm is greedy if querying variables in descending order of weight. Intuitively, any greedy query algorithm computing $f$ should be the most efficient approach to estimate the sum $w_{0}+\sum_{i=1}^{n}w_{i}x_{i}$ and determine whether the sum is over zero or not. The next proposition shows this intuition is indeed correct.
Proposition B.17.
The greedy query algorithm is optimal for every LTF in terms of $\mathrm{D}_{\mathrm{ave}}(f)$.
Proof.
We prove this by induction on $\mathrm{D}(f)$. Assume $w_{1}\geq\cdots\geq w_{n}$ are non-negative without loss of generality. Observe that every LTF $f=\mathrm{sgn}(w_{0}+\sum_{i=1}^{n}w_{i}x_{i})$ is constant if and only if $w_{0}\geq\sum w_{i}$ or $w_{0}<-\sum w_{i}$.
Base case ($\mathrm{D}(f)=1$):
A zero-error decision tree that queries only the variable $x_{i}$ computes $f$ if and only if $w_{0}+w_{i}\geq\sum_{j\neq i}w_{j}$ and $w_{0}-w_{i}<-\sum_{j\neq i}w_{j}$. If we query the variable $x_{1}$ instead, given that $w_{1}\geq w_{i}$ for all $i=1,2,\ldots,n$, the following inequalities still hold:
  • $w_{0}+(w_{1}-w_{i})\geq w_{0}+(w_{i}-w_{1})\geq\sum_{j\neq 1,i}w_{j}\Longrightarrow w_{0}+w_{1}\geq\sum_{j\neq 1}w_{j}$.
  • $w_{0}-(w_{1}-w_{i})\leq w_{0}-(w_{i}-w_{1})<-\sum_{j\neq 1,i}w_{j}\Longrightarrow w_{0}-w_{1}<-\sum_{j\neq 1}w_{j}$.
Inductive step ($\mathrm{D}(f)>1$):
Figure 2: Inductive step (Proposition B.17)
Consider any optimal decision tree $T$ that computes $f$ with root labeled by some integer $i\neq 1$. The tree $T$ has two subtrees $\Gamma,\Delta$, which compute the subfunctions $f|_{x_{i}=+1}$ and $f|_{x_{i}=-1}$, respectively. The trees $\Gamma,\Delta$ further branch into $\alpha,\beta,\gamma,\delta$. By the induction hypothesis, $\Gamma,\Delta,\alpha,\beta,\gamma,\delta$ are greedy trees, and the roots of $\Gamma,\Delta$ are labels by $1$.
If neither $f|_{x_{i}=+1}$ nor $f|_{x_{i}=-1}$ is constant, then both $\Gamma,\Delta$ are not empty. We can obtain a greedy tree computing $f$ with the same average depth by properly relabeling the internal nodes and rearranging the subtrees $\alpha,\beta,\gamma,\delta$. Refer to Figure 2 for illustration.
Suppose either $f|_{x_{i}=-1}$ or $f|_{x_{i}=+1}$ is constant, and assume $f|_{x_{i}=-1}$ is the constant one without loss of generality. If $f|_{x_{i}=-1}\equiv+1$, then $f|_{x_{i}=+1}\equiv+1$ (since $w_{1},\ldots,w_{n}$ are non-negative, and $f$ is monotone). Thus, we have $f\equiv+1$, contradicting the assumption that $\mathrm{D}(f)>1$. Consequently, it must be that $f|_{x_{i}=-1}\equiv-1$, leading to $w_{0}-w_{i}<-\sum_{j\neq i}w_{j}$. Given that $w_{1}\geq w_{i}$ for all $i=1,2,\ldots,n$, it follows that
$w_{0}+(w_{i}-w_{1})\leq w_{0}+(w_{1}-w_{i})<-\sum_{j\neq 1,i}w_{j},$
implying $f|_{x_{i}=+1,x_{1}=-1}\equiv-1$. Therefore, $\beta,\Delta$ are empty (i.e., they have no internal nodes), and $f|_{x_{1}=+1,x_{i}=+1}$ is the only non-constant subfunction. Since querying proceeds only if $x_{1}=x_{i}=+1$, a decision tree querying $x_{1}$ before $x_{i}$ and applying the subtree $\alpha$ to the subfunction $f|_{x_{1}=+1,x_{i}=+1}$ also computes $f$ and is greedy and optimal. ∎
B.5 Tribes function
Definition B.18.
The tribes functions are defined by $\mathrm{Tribes}_{w,s}(x)=\bigvee_{i=1}^{s}\bigwedge_{j=1}^{w}x_{ij}.$
The tribes functions serve as a common example in boolean functions analysis. We compute the exact value of $\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})$.
Proposition B.19.
$\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})=2(2^{w}-1)\left(1-(1-\frac{1}{2^{w}})^{s}\right).$
Proof.
Let $\gamma=\Pr_{{x}\in\{0,1\}^{n}}[\mathrm{Tribes}_{w,s}({x})=0]=(1-\frac{1}{2^{w}})^{s}$. On one hand, one can easily verify that $\mathrm{{Var}}[\mathrm{Tribes}_{w,s}]=4\gamma(1-\gamma)$ and $\mathrm{{Inf}}_{i}[\mathrm{Tribes}_{w,s}]=\Pr_{{x}\in\{0,1\}^{n}}[\mathrm{Tribes}_{w,s}(x)\neq\mathrm{Tribes}_{w,s}({x}^{\oplus i})]=\frac{2\gamma}{2^{w}-1}$ for any $i=1,2,\ldots,n$. Thus, we have $\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})\geq 2(2^{w}-1)(1-(1-\frac{1}{2^{w}})^{s})$ by the OSSS inequality (Theorem A.1). On the other hand, we prove
$\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})\leq 2(2^{w}-1)\left(1-(1-\frac{1}{2^{w}})^{s}\right)$ (22)
by induction on $s$. By Corollary B.2, $\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,1})=\mathrm{D}_{\mathrm{ave}}(\mathrm{OR}_{w})=2(1-\frac{1}{2^{w}})$, which proves (22) for $s=1$. Our strategy is to query the clauses one by one until the output of that clause is determined; if the clause evaluates to 1, then the output can be determined. Thus, by the induction hypothesis, we have
$\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})$ $\leq\left(\sum_{k=1}^{w}\frac{1}{2^{k}}\left(k+\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s-1})\right)\right)+\frac{w}{2^{w}}$
$=2\left(1-\frac{1}{2^{w}}\right)+\left(1-\frac{1}{2^{w}}\right)\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s-1})$
$=2(2^{w}-1)\left(1-(1-\frac{1}{2^{w}})^{s}\right),$
completing the induction proof of (22). ∎
The following propositions give the exact value of the average sensitivity, the average block sensitivity and the average certificate complexity of the tribes functions.
Proposition B.20.
[Proposition 4.13 in [ODo14]] $\mathrm{s}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})=\frac{2ws}{2^{w}-1}\left(1-\frac{1}{2^{w}}\right)^{s}$.
Proposition B.21.
$\mathrm{bs}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})=\mathrm{C}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})=\left(1-\left(1-\frac{1}{2^{w}}\right)^{s}\right)w+\left(1-\frac{1}{2^{w}}\right)^{s}s.$
Proof.
Let $f=\mathrm{Tribes}_{w,s}$. Observe that $\mathrm{Tribes}_{w,s}(x)=1$ if and only if $x_{i,1}=\cdots=x_{i,w}=1$ for some integer $i\in\{1,\ldots,s\}$, and that $\mathrm{Tribes}_{w,s}(x)=0$ if and only if $x_{1,j_{1}}=\cdots=x_{s,j_{s}}=0$ for some integers $j_{1},\cdots,j_{s}\in\{1,\ldots,w\}$. Thus, each input $x$ has a $1$-certificate of size $w$ or a 00-certificates of size $s$, which implies that
$\mathrm{C}_{\mathrm{ave}}(f)$ $=\Pr_{x\in\{0,1\}^{n}}[f(x)=1]\cdot w+\Pr_{x\in\{0,1\}^{n}}[f(x)=0]\cdot s$
$=\left(1-\left(1-\frac{1}{2^{w}}\right)^{s}\right)w+\left(1-\frac{1}{2^{w}}\right)^{s}s.$
For notational convenience, let $\mathrm{Tribes}_{w,s}(x)=C_{1}\vee C_{2}\vee\cdots\vee C_{s}$, where $C_{i}=\bigwedge_{j=1}^{w}x_{ij}$. If $\mathrm{Tribes}_{w,s}(x)=1$, in which clauses $C_{i_{1}},C_{i_{2}},\ldots,C_{i_{k}}$ evaluate to $1$, it is evident that there exist $w$ (but no more) disjoint blocks
$B_{1}=\{x_{i_{1},1},\cdots,x_{i_{k},1}\},\cdots,B_{w}=\{x_{i_{1},w},\cdots,x_{i_{k},w}\}$
such that $f(x)\neq f(x^{\oplus B_{i}})$ for any $1\leq i\leq w$. If $\mathrm{Tribes}_{w,s}(x)=0$, there are $s$ (but no more) disjoint blocks
$B_{i}=\{x_{i,j}:x_{i,j}=0\text{ for }j=1,2,\ldots,w\},$
where $i=1,2,\ldots,s$, such that $f(x)\neq f(x^{\oplus B_{i}})$. Thus, in both cases, $\mathrm{bs}_{x}(f)$ equals $\mathrm{C}_{x}(f)$, which implies that $\mathrm{bs}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})=\mathrm{C}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})$. ∎
Corollary B.22.
$\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})=\Theta(s)$ (respectively, $\mathrm{R}_{\mathrm{ave}},\mathrm{C}_{\mathrm{ave}},\mathrm{bs}_{\mathrm{ave}}$) when $s<2^{w}\ln 4$.
Proof.
Let $f=\mathrm{Tribes}_{w,s}$. By Proposition B.21, $\mathrm{bs}_{\mathrm{ave}}(f)=\mathrm{C}_{\mathrm{ave}}(f)$. By Proposition 3.1, 3.2, we have $\mathrm{D}_{\mathrm{ave}}(f)\geq\mathrm{R}_{\mathrm{ave}}(f)=\Omega(\mathrm{bs}_{\mathrm{ave}}(f))$.
By Proposition B.19, $\mathrm{D}_{\mathrm{ave}}(f)=2(2^{w}-1)(1-(1-\frac{1}{2^{w}})^{s})\leq 2(2^{w}-1)(1-(1-\frac{s}{2^{w}}))\leq 2s=O(s)$. By Proposition B.21,
$\mathrm{bs}_{\mathrm{ave}}(f)$ $\geq s\cdot\left(1-\frac{1}{2^{w}}\right)^{s}$
$\geq s\cdot\left(1-\frac{1}{2^{w}}\right)^{2^{w}\ln 4}$ $\bigg{(}s<2^{w}\ln 4\bigg{)}$
$\geq s\cdot e^{-\ln 4}\left(1-\frac{\ln 4}{2^{w}}\right).$ $\left(\left(1+\frac{x}{n}\right)^{n}\geq e^{x}\left(1-\frac{x^{2}}{n}\right)\text{ for }n\geq 1,x\leq|n|\right)$
Thus, we conclude that $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$, $\mathrm{C}_{\mathrm{ave}}(f)$, $\mathrm{bs}_{\mathrm{ave}}(f)$ are all $\Theta(s)$. ∎
To lower bound $\mathrm{R}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})$ when $s\geq 2^{w}\ln 4$, we use the technique of Ambainis and de Wolf, which was used to prove the exponential gap between $\mathrm{R}_{\mathrm{ave}}(f)$ and $\mathrm{Q}_{\mathrm{ave}}(f)$. For completeness, we reproduce their proof here.
Let $\mathcal{D}_{0}$ (respectively, $\mathcal{D}_{1}$) be the uniform distribution over inputs $x$ such that $f(x)=0$ (respectively, $f(x)=1$). Let $\mathcal{U}$ be the uniform input distribution over $\{0,1\}^{n}$. Say a randomized decision tree $\mathcal{T}$ distinguishes between $\mathcal{D}_{0}$ and $\mathcal{U}$ if $\Pr_{x\sim\mathcal{D}_{0},T\sim\mathcal{T}}[T(x)=1]<1/3$ and $\Pr_{x\sim\mathcal{U},T\sim\mathcal{T}}[T(x)=0]<1/3$. Note that $\mathcal{T}$ does not necessarily compute $f$.
Lemma B.23 (Lemma 4.3 in [AD01]).
Suppose $\Pr_{x\sim\mathcal{U}}[f(x)=0]\leq 1/4$. There exists a randomized decision tree $\mathcal{T}$ that distinguishes $\mathcal{D}_{0}$ and $\mathcal{U}$, and always uses $O(\mathrm{R}_{\mathrm{ave}}(f))$ queries on all inputs.
Proof.
Let $p=\Pr_{x\sim\mathcal{U}}[f(x)=0]$. Let $\mathcal{A}$ be an optimal bounded-error randomized decision tree computing $f$ such that $m=\mathrm{R}_{\mathrm{ave}}(f)=\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}[\mathrm{cost}(\mathcal{A},x)]$. Without loss of generality, assume $\mathcal{A}$ computes $f$ with error $q=\frac{1}{6}-\frac{p}{2}$. Since $p\leq\frac{1}{4}$, we have $q\geq\frac{1}{24}$. We construct a (randomized) query algorithm $\mathcal{T}$ that distinguishes $\mathcal{D}_{0}$ and $\mathcal{U}$ as follows:
  1. 1.
    Run $\mathcal{A}$ until it stops or queries $\frac{m}{q}$ variables.
  2. 2.
    Output the answer of $\mathcal{A}$ if it stops; otherwise, output 00.
Obviously, on every input, the algorithm $\mathcal{T}$ queries at most $\frac{m}{q}\leq 24m=O(m)$ variables.
From the construction of $\mathcal{T}$, we have
$\Pr_{x\sim\mathcal{U},T\sim\mathcal{A}}\left[T(x)=0\right]$ $\leq\Pr_{x\sim\mathcal{U}}[f(x)=0]+\Pr_{x\sim\mathcal{U}}[f(x)=1]\cdot\Pr_{x\sim\mathcal{U},T\sim\mathcal{A}}\left[T(x)=0\mid f(x)=1\right]$
$\leq p+(1-p)\cdot q.$
By Markov’s inequality, over a random input from $\mathcal{U}$, the probability that $\mathcal{A}$ queries more than $\frac{m}{q}$ variables is at most by $q$. Thus, we have
$\Pr_{x\sim\mathcal{U},T\sim\mathcal{T}}[T(x)=0]$ $\leq\Pr_{x\sim\mathcal{U},T\sim\mathcal{A}}[T(x)=0]+\Pr_{x\sim\mathcal{U},T\sim\mathcal{A}}\left[\mathrm{cost}(T,x)>\frac{m}{q}\right]$
$\leq p+(1-p)\cdot q+q<p+2q=\frac{1}{3}.$
On the other hand, over the uniform random input from $\mathcal{D}_{0}$, $\mathcal{T}$ outputs $1$ with probability $q\leq\frac{1}{6}$, since $f(x)=0$ for any input $x$ drawn from $\mathcal{D}_{0}$. Therefore, we conclude that $\mathcal{T}$ distinguish between $\mathcal{D}_{0}$ and $\mathcal{U}$. ∎
Proposition B.24.
$\mathrm{D}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})=\Theta(2^{w})$ and $\mathrm{R}_{\mathrm{ave}}(\mathrm{Tribes}_{w,s})=\Theta(2^{w})$ when $s\geq 2^{w}\ln 4$.
Proof.
Let $f=\mathrm{Tribes}_{w,s}$. By Proposition B.19, $\mathrm{D}_{\mathrm{ave}}(f)=2(2^{w}-1)(1-(1-\frac{1}{2^{w}})^{s})\leq 2\cdot 2^{w}$. By Proposition 3.1, we have $\mathrm{R}_{\mathrm{ave}}(f)\leq\mathrm{D}_{\mathrm{ave}}(f)=O(2^{w})$. We shall prove that $\mathrm{R}_{\mathrm{ave}}(f)=\Omega(2^{w})$ when $s\geq 2^{w}\ln 4$.
Assume for contradiction that $\mathrm{R}_{\mathrm{ave}}(f)=o(2^{w})$. Note that
$\Pr_{x\in\{0,1\}^{n}}\left[f(x)=0\right]=\left(1-\frac{1}{2^{w}}\right)^{s}\leq\left(1-\frac{1}{2^{w}}\right)^{2^{w}\ln 4}\leq e^{-\ln 4}=\frac{1}{4}.$
By Lemma B.23, there exists a randomized decision tree $\mathcal{T}$ distinguishing between $\mathcal{D}_{0}$ and $\mathcal{U}$, and using $o(2^{w})$ queries on all inputs, i.e., $\mathrm{cost}(T,x)=o(2^{w})$ for any input $x\in\{0,1\}^{n}$ and any decision tree $T$ drawn from $\mathcal{T}$.
For any decision tree $T$ querying at most $t=t(T)$ variables, $T$ queries no more than $t$ clauses; without loss of generality, assume that every decision tree $T$ drawn from $\mathcal{T}$ queries only the first $t=o(2^{w})$ clauses $C_{1},\ldots,C_{t}$, where $C_{i}=\bigwedge_{j=1}^{w}x_{i,j}$ for $i=1,2,\ldots,t$. Despite that $T$ could always compute $f(x)$ on those inputs $x$ satisfying $\bigvee_{i=1}^{t}C_{i}=1$, $T$ gets such inputs with probability
$\Pr_{x\sim\mathcal{U}}\left[\bigvee_{i=1}^{t}C_{i}=1\right]=1-\left(1-\frac{1}{2^{w}}\right)^{t}=o(1).$
[…]
$\sum_{P\in\mathcal{P}}\Pr_{x\in\mathcal{D}}\left[P\text{ accepts }x\right]=1$ (23)
for any input distribution $\mathcal{D}$.
Let $\mathcal{V}$ be the uniform distribution over inputs $x\in\{0,1\}^{n}$ satisfying $\bigvee_{i=1}^{t}C_{i}=0$. Write input $x=(y,z)$, where $y\in\{0,1\}^{tw}$ and $z\in\{0,1\}^{n-tw}$. The distribution $\mathcal{D}_{0}$ can be viewed as a distribution that samples $y$ from the set $Y=\{y\in\{0,1\}^{tw}:\bigvee_{i=1}^{t}C_{i}=0\}$ and $z$ from the set $Z_{0}=\{z\in\{0,1\}^{n-tw}:\bigvee_{i=t+1}^{n}C_{i}=0\}$ independently. The distribution $\mathcal{V}$ can be viewed as a distribution that samples $y$ from the same set $Y$, and samples $z$ from the set $Z=\{0,1\}^{n-tw}$. Since $T$ queries only the first $t$ clauses (and thus at most the first $tw$ variables), we have
$\Pr_{x\sim\mathcal{D}_{0}}[P\text{ accepts }x]=\Pr_{y\in Y}\left[P\text{ accepts }y\right]=\Pr_{x\sim\mathcal{V}}[P\text{ accepts }x]$ (24)
for any $P\in\mathcal{P}$.
On any input $x$ accepted by $P$, the output of the tree $T$ is constant, i.e., $T(x)\equiv b$, no matter what distribution $x$ comes from. In other words, for any input distribution $\mathcal{D}$,
$\Pr_{x\sim\mathcal{D}}\left[T(x)=b\mid P\text{ accepts }x\right]=1.$
Therefore, we have
$\ \Pr_{x\sim\mathcal{D}_{0}}\left[T(x)=0\mid P\text{ accepts }x\right]+\Pr_{x\sim\mathcal{V}}\left[T(x)=1\mid P\text{ accepts }x\right].$
$=$ $\ 1+\Pr_{x\sim\mathcal{D}_{0}}\left[T(x)=0\mid P\text{ accepts }x\right]-\Pr_{x\sim\mathcal{V}}\left[T(x)=0\mid P\text{ accepts }x\right]$
$=$ $\ 1$ (25)
Combining (23), (24), (25), we have
$\ \Pr_{x\sim\mathcal{D}_{0}}[T(x)=0]+\Pr_{x\sim\mathcal{V}}[T(x)=1]$
$=$ $\ \left(\sum_{P\in\mathcal{P}}\Pr_{x\sim\mathcal{D}_{0}}[P\text{ accepts }x]\cdot\Pr_{x\sim\mathcal{D}_{0}}\left[T(x)=0\mid P\text{ accepts }x\right]\right)$
$\ +\left(\sum_{P\in\mathcal{P}}\Pr_{x\sim\mathcal{V}}[P\text{ accepts }x]\cdot\Pr_{x\sim\mathcal{V}}\left[T(x)=1\mid P\text{ accepts }x\right]\right)$
$=$ $\ \sum_{P\in\mathcal{P}}\Pr_{x\sim\mathcal{V}}[P\text{ accepts }x]\left(\Pr_{x\sim\mathcal{D}_{0}}\left[T(x)=0\mid P\text{ accepts }x\right]+\Pr_{x\sim\mathcal{V}}\left[T(x)=1\mid P\text{ accepts }x\right]\right)$
$=$ $\ \sum_{P\in\mathcal{P}}\Pr_{x\sim\mathcal{V}}[P\text{ accepts }x]$
$=$ $\ 1$ (26)
Since
$\Pr_{x\sim\mathcal{U}}\left[T(x)=1\right]\leq\Pr_{x\sim\mathcal{V}}\left[T(x)=1\right]+\Pr_{x\sim\mathcal{U}}\left[\bigvee_{i=1}^{t}C_{i}=1\right]=\Pr_{x\sim\mathcal{V}}\left[T(x)=1\right]+o(1),$
we have $\Pr_{x\sim\mathcal{D}_{0}}[T(x)=0]+\Pr_{x\sim\mathcal{U}}[T(x)=1]\leq 1+o(1)$ by (26), which implies either $\Pr_{x\sim\mathcal{D}_{0}}[T(x)=0]\leq\frac{1}{2}+o(1)$ or $\Pr_{x\sim\mathcal{U}}[T(x)=1]\leq\frac{1}{2}+o(1)$. Recall that $T$ is a decision tree drawn from $\mathcal{T}$ arbitrarily. Therefore, we conclude that either $\Pr_{x\sim\mathcal{D}_{0},T\sim\mathcal{T}}[T(x)=1]\geq\frac{1}{2}-o(1)$ or $\Pr_{x\sim\mathcal{U},T\sim\mathcal{T}}[T(x)=0]\geq\frac{1}{2}-o(1)$. Thus, $\mathcal{T}$ cannot distinguish between $\mathcal{D}_{0}$ and $\mathcal{U}$. Contradiction. ∎